### v1

# Reinforcement Inference: Leveraging Uncertainty for Self-Correcting Language Model Reasoning

Abstract
Modern large language models (LLMs) are often evaluated and deployed under a one-shot, greedy inference protocol, especially in professional settings that require deterministic behavior. This regime can systematically under-estimate a fixed model’s true capability: many errors arise not from missing knowledge, but from premature commitment under internal ambiguity. We introduce Reinforcement Inference, an entropy-aware inference-time control strategy that uses the model’s own uncertainty to selectively invoke a second, more deliberate reasoning attempt, enabling stronger performance without any retraining.
On 12,032 MMLU-Pro questions across 14 subjects, using DeepSeek-v3.2 with deterministic decoding in a zero-shot setting, Reinforcement Inference improves accuracy from 60.72% to 84.03%, while only incurring 61.06% additional inference calls. A 100% re-asking ablation reaches 84.35%, indicating that uncertainty-aware selection captures most of the attainable improvement with substantially less compute. Moreover, a prompt-only ablation underperforms the baseline, suggesting that the gains are not explained by generic “ your output had high entropy, think step-by-step” prompting alone.
Beyond providing a practical inference-time upgrade, our results suggest a broader entropy-aware paradigm for measuring and expanding model capability: because modern decoder-based models generate outputs autoregressively, entropy and related confidence measures arise naturally as first-class control signals during generation. The resulting gap between one-pass greedy inference and uncertainty-conditioned deliberation offers a diagnostic lens on an LLM’s latent reasoning horizon and motivates future training objectives that explicitly constrain correctness–confidence alignment.
I Introduction
Despite the remarkable performance, LLMs can produce confidently incorrect answers. A well-known challenge is that models often lack calibrated confidence – they may not recognize when their answers are likely wrong [1, 2]. This can lead to misleading outputs in high-stakes settings (for example, an LLM hallucinating a plausible-sounding but false explanation with full certainty [3, 4]). If an AI could know when it is unsure, it might mitigate errors by double-checking or deferring answers. Indeed, prior studies on tasks like multiple-choice QA have noted that an LLM’s softmax probabilities carry information about correctness [5, 2]. Intuitively, a model that assigns one option nearly 100% probability is likely confident (and often correct), whereas if it spreads probability roughly uniformly across all options, it is essentially guessing [5, 2]. Recent evaluations confirm that wrong answers tend to have lower maximum softmax probability (MSP) than correct answers in well-trained models [2, 6]. In short, uncertainty correlates with mistakes.
How can we use this insight to improve model reasoning? One approach is to have the model abstain or refuse to answer when uncertainty is high [7]. Refusal might be safer in open-ended applications, but in many scenarios we do want the model to attempt an answer – just with more care if it was initially unsure. Humans do this naturally: if unsure of an answer, a person might pause and double-check their reasoning or try a different approach before responding. This inspires the question: Can an LLM similarly “think twice” when it hesitates, to produce a more accurate answer, all at inference time?
Figure 1: System structure of Reinforcement Inference.
In this paper, we propose an inference-time framework called Reinforcement Inference to realize this idea. The core concept is simple: the model’s own inferred uncertainty and MSP are used as a trigger for a second reasoning attempt. If the model’s first answer had high entropy – indicating confusion or low confidence – we prompt the model again with a special instruction that it was uncertain and should reason more thoroughly. Crucially, we do not reveal the correct answer or provide any external feedback; the model must rely on its existing knowledge, but is now explicitly guided to be careful. This procedure requires no additional training or parameter updates. It functions as a form of test-time control or meta-cognition for the model, akin to a student checking their work upon sensing uncertainty. As we will show, even this light-touch intervention can yield significant accuracy improvements on challenging questions that the model initially struggled with.
Before diving into technical details, we emphasize the broader motivation of this work: modern LLM capability is often not fully realized by the default inference protocol. In many deployed settings—especially in professional and industrial workflows that require deterministic behavior—models are typically used with greedy decoding and a single forward pass. This “one-shot, greedy” regime can systematically under-estimate what a fixed model is actually capable of: many failures are not due to missing knowledge, but due to shallow first attempts, premature commitment, or unresolved internal ambiguity. Reinforcement Inference provides a practical way to probe and unlock a model’s latent capability horizon without any retraining. The large accuracy jump we observe is therefore not merely a reprompting trick; it is evidence that a substantial portion of the model’s reasoning capacity is available at inference time but remains untapped under the standard one-pass protocol.
Methodologically, our key idea is to treat uncertainty as a first-class control signal during inference. Entropy and MSP expose when the model is effectively “guessing” under its own posterior, and we convert this signal into an adaptive compute policy: allocate a second reasoning pass only when the model itself indicates hesitation. This yields an efficiency–accuracy paradigm suited to real systems: predictable latency on easy cases, and deliberate reasoning only when needed.
More broadly, we argue that entropy-aware inference suggests an entropy-aware training direction. For professional models where correctness and determinism matter, training objectives and decoding policies should explicitly constrain the relationship between correctness and confidence (e.g., penalizing high-confidence errors and encouraging calibrated decisiveness). Our work offers a concrete blueprint for such future designs: (i) use entropy/MSP to define when additional reasoning is warranted, (ii) quantify the resulting “capability gap” between one-pass greedy inference and uncertainty-conditioned deliberation, and (iii) use this gap as a diagnostic signal for both decoding and model optimization. In this sense, Reinforcement Inference is a step toward a broader paradigm in which uncertainty is not only measured, but also systematically acted upon.
II Contributions
Our contributions can be summarized as follows:
  • A “capability-horizon” view of inference without retraining: We argue and empirically demonstrate that the default one-shot, greedy inference protocol can materially under-estimate what a fixed LLM can do. Reinforcement Inference provides a practical way to unlock this latent reasoning horizon without any retraining, which is directly relevant to industrial deployments where model weights cannot be updated frequently.
  • Entropy-aware inference as a systems paradigm: We introduce an entropy-aware inference-time control strategy that treats uncertainty as a first-class signal for allocating extra compute. This yields a deployment-friendly accuracy–latency trade-off: easy (confident) cases remain single-pass and deterministic, while only uncertain cases trigger an additional deliberate attempt.
  • Evidence that uncertainty is predictive (first-round trial, no leakage): On 12,032 MMLU-Pro questions, we report a clear separation between correct and incorrect first-pass answers (mean entropy 0.845 vs. 1.567 nats; mean MSP 0.554 vs. 0.249). Importantly, in this first-round trial the model never sees any ground-truth labels; correctness is used only for post-hoc evaluation, so there is no ground-truth leakage to the model.
  • Experimental validation + 16-run threshold sweep: We validate large gains from uncertainty-conditioned re-asking at scale, and we additionally run a 16-configuration sweep of $(\tau_{H},\tau_{\mathrm{MSP}})$ to study robustness and identify compute-efficient operating points for future validation. Across the sweep, final accuracy remains stable (about 83.69%–84.34%) while re-ask rates vary (52.80%–76.31%), demonstrating that the approach is not brittle and enabling practitioners to choose thresholds based on budget.
  • A practical inference-parameter paradigm: Beyond a single fixed configuration, we frame uncertainty-conditioned re-asking as a tunable inference policy with meaningful “best” operating regions. In practice, a model can have near-optimal threshold ranges where investing a controlled amount of extra compute yields substantially higher correctness on hard instances, while preserving deterministic behavior and avoiding unnecessary overhead on easy ones.
III Related Work
III-A Confidence and Uncertainty in LLMs
There is growing interest in measuring how well an LLM’s predicted probabilities reflect its actual likelihood of being correct [2, 15], including recent work on whether LLMs can reason explicitly about their own confidence [44]. Maximum softmax probability (MSP) has been used as a simple proxy for model confidence, with studies showing that lower MSP correlates with wrong answers and even with out-of-distribution inputs in both vision and NLP models [2, 15]. For instance, Plaut et al. (2024) provide strong evidence that in multiple-choice QA, an LLM’s softmax distribution can predict whether its answer is correct [16, 17]. They leveraged this by allowing the model to abstain on low-confidence answers, improving overall accuracy by avoiding some mistakes [46, 18, 17]. Other works explore more complex uncertainty quantification methods (Bayesian approaches, ensembles, or probing internal states [17, 19]) and calibrating models via fine-tuning to output better confidence estimates [20, 21]. Our work aligns with these efforts in that we use the model’s own probability distribution as an uncertainty signal. However, instead of training the model to be calibrated or to abstain, we use uncertainty to trigger a self-correction process at inference time.
III-B Chain-of-Thought and Self-Consistency
Prompting strategies like chain-of-thought (CoT) encourage models to generate intermediate reasoning steps, which often boosts performance on complex tasks [8]. Building on CoT, the self-consistency approach samples multiple independent reasoning paths and then selects the most common answer among them [8, 9]. Wang et al. (2023) showed that self-consistency significantly improves accuracy on math and commonsense questions by effectively ensembling reasoning outcomes [9]. While powerful, self-consistency incurs substantial computational cost (many model forward passes per query) and always uses multiple attempts even if the model was confident on the first try. In contrast, our method performs a second pass selectively – only for cases where the first pass was uncertain. Thus, Reinforcement Inference can be seen as a targeted, single-step variant of reasoning augmentation. It resembles asking the model to “try a different line of reasoning” but only when warranted by low confidence, saving computation on easier queries.
III-C Verifier Models and Answer Re-Ranking
Another line of work improves answer correctness by introducing a verification stage. One paradigm is to train a separate verifier classifier that judges whether the model’s answer is correct (sometimes using explanations or retrieved evidence), and then use that to accept or reject answers [10, 11]. Other approaches prompt the model itself to check its answer. For example, Chain-of-Verification (CoVe) has the model generate an initial answer, then pose follow-up verification questions to itself, answer those, and finally produce a revised answer [22, 23]. This multi-step deliberation can reduce factual hallucinations [24, 25]. Similarly, self-critique techniques have the model reflect on and potentially correct its first answer by analyzing its own output for errors [26, 27]. These verification-based methods improve reliability, but they often rely on complex prompt programs or fine-tuned models for critique. Our approach is much simpler: we do not generate explicit verification questions or use a separate model; we just give a single additional prompt reminding the model to be careful. In effect, Reinforcement Inference is a lightweight self-reflection step triggered by uncertainty, rather than a full-fledged self-dialogue or learned verifier.
III-D Supervised Retraining for Reasoning
A straightforward way to improve a model’s accuracy or calibration is additional training. Fine-tuning on high-quality reasoning data (e.g. with CoT exemplars or human explanations) can make a model more reliable on tricky problems. Reinforcement Learning from Human Feedback (RLHF) is also widely used to align models to preferred answers and avoid obvious mistakes. For example, recent work has fine-tuned LLMs specifically to output better calibrated uncertainty estimates or to know when to say “I don’t know” [13, 28]. However, these methods require collecting labeled data or rewards and then retraining the model – which may not be feasible for very large or closed-source models. In contrast, our method introduces no new training at all: it works as an inference-time plug-in that could be applied to any pre-trained model (even black-box API models, provided one can extract a confidence measure like probabilities or logits).
IV Methodology
IV-A Uncertainty Metrics for Model Answers
We focus on multiple-choice questions, where the model must choose one correct option from a given set (e.g. A, B, C, D, E, F, G, H, I, J). When prompted with a question and answer choices, the model produces a probability distribution $P$ over the options.
In the main experiments, we compute entropy in nats. For cross-dataset and cross-model evaluation, we report normalized entropy to ensure scale invariance across tasks with different numbers of answer options.
We define two metrics from this distribution:
  • Entropy ($H$): This measures the overall uncertainty or spread of the distribution. Formally,
    $H(P)=-\sum_{i=1}^{K}P_{i}\ln P_{i},$
    where $K$ is the number of options ($K=10$ in MMLU-Pro). We express entropy in nats; in this case, a uniform distribution has $H=\ln K$ (e.g., $H=\ln 10\approx 2.303$ when $K=10$), while a peaked distribution (e.g. one option with $P\approx 1$ and the others near 0) has entropy close to 0. In our analysis, we compute $H$ using the model’s normalized probabilities over the provided choices. (In the extreme deterministic case, where $p_{\max}=1$, we have $H=-\log p_{\max}$.)
  • Maximum Softmax Probability ($MSP$): This is the model’s confidence in its top choice, defined as
    $\mathrm{MSP}(P)=\max_{1\leq i\leq K}P_{i}.$
    For example, if the model selects option A with probability 0.7, then $\mathrm{MSP}=0.7$. A higher $\mathrm{MSP}$ indicates the model was more confident in the answer it chose, whereas a low $\mathrm{MSP}$ (e.g. 0.25) indicates it was essentially guessing among several options.
  • Normalized Entropy ($\bar{H}$): Because $H$ scales with $K$, we also consider a normalized form
    $\bar{H}(P)=\frac{H_{\mathrm{nat}}(P)}{\ln K}\in[0,1],$
    which enables comparing uncertainty across tasks with different numbers of options. This normalization is a monotonic rescaling and does not alter the triggering behavior.
These metrics are inversely related (a high $\mathrm{MSP}$ usually implies low entropy, and vice versa), but we consider both for completeness. Entropy captures the notion of hesitation in the distribution, while $\mathrm{MSP}$ directly captures confidence in the chosen answer. From a Bayesian decision perspective, if the model’s posterior over options is $P$, the Bayes-optimal 0–1 decision is $\hat{y}=\arg\max_{i}P_{i}$, and $1-\mathrm{MSP}(P)$ upper-bounds the posterior risk of choosing $\hat{y}$ under the model’s own beliefs. Motivated by calibration theory, we also define two proper scoring rules [45], which we report as secondary diagnostics in Section VI-A:
  • Brier score: for one-hot label $y$,
    $\mathrm{BS}(P,y)=\sum_{i=1}^{K}(P_{i}-\mathbb{I}[i=y])^{2},$
    which penalizes both overconfidence and underconfidence.
  • Negative log-likelihood (NLL):
    $\mathrm{NLL}(P,y)=-\log P_{y},$
    which corresponds to the cross-entropy loss of the true option.
We hypothesize that correct answers will on average have lower entropy and higher $\mathrm{MSP}$ than incorrect ones – indicating the model tends to be more decisive when it is right, and more ambivalent when it is wrong.
To test this hypothesis, we ran DeepSeek-v3.2 on a large set of multiple-choice questions and extracted the model’s answer distribution over the provided options from token log-probabilities (converted to normalized probabilities). We then computed $H$ and Max $P$ for each question and compared these uncertainty metrics between items that were answered correctly vs. incorrectly post hoc.
This analysis is inspired by standard calibration-style evaluation practice: rather than claiming any specific distributional assumptions, we report descriptive statistics and visualize the separation between the two groups.
IV-B Empirical Trigger Criterion
We use uncertainty to decide whether to invoke a second reasoning attempt. We compute entropy $H(P)$ using the natural logarithm, yielding values in nats. For ten-way multiple-choice questions (MMLU-Pro), the maximum entropy is $\ln 10\approx 2.303$ nats.
Main setting (MMLU-Pro, 10-way). To avoid any threshold selection leakage, our main setting uses fixed, a priori thresholds motivated by the interpretation of entropy/MSP under a Bayesian 0–1 decision framework. For $K=10$ options, the maximum entropy is $\ln K\approx 2.303$ nats.
We adopt an entropy threshold of $\tau_{H}=0.8$ nats, corresponding to roughly 35% of the maximum uncertainty,
$\frac{\tau_{H}}{\ln 10}\approx 0.347,$
which selects predictions with a clearly peaked but not fully deterministic posterior. In parallel, we require $\mathrm{MSP}(P)\geq\tau_{\mathrm{MSP}}=0.6$, which upper-bounds the model’s posterior risk by $1-\tau_{\mathrm{MSP}}=0.4$.
Together, these thresholds identify predictions that are both locally confident in the chosen answer and globally low in distributional ambiguity, while avoiding overly conservative or near-deterministic regimes.
We trigger re-asking when either criterion indicates hesitation:
$\text{trigger}=\mathbb{I}[H(P)>\tau_{H}\ \lor\ \mathrm{MSP}(P)<\tau_{\mathrm{MSP}}].$
Cross-dataset transfer (MMLU, 4-way). When transferring the same normalized-entropy threshold to 4-way multiple-choice tasks, we use the equivalent unnormalized entropy threshold
$\tau_{H}^{(K=4)}=\bar{H}_{\tau}\ln 4\approx 0.78\ \text{nats},$
while keeping $\tau_{\mathrm{MSP}}=0.4$ fixed.
Although these thresholds were guided by an initial trial analysis, we emphasize that selecting $\tau_{H},\tau_{\mathrm{MSP}}$ after inspecting performance can introduce optimistic bias. We discuss this “privileged” threshold selection issue and its implications in Section VI-F and Limitations.
The Reinforcement Inference procedure can be summarized as Algorithm 1:
  1. 1.
    First Pass: Present the question to the model in the standard zero-shot format and obtain its answer (along with the probability distribution $P$ over answer options).
  2. 2.
    Uncertainty Check: Compute entropy $H(P)$ and maximum softmax probability $\mathrm{MSP}(P)$. If $H(P)\leq\tau_{H}$ and $\mathrm{MSP}(P)\geq\tau_{\mathrm{MSP}}$, consider the model confident and accept its first answer as final.
  3. 3.
    Reinforcement Prompt (Second Pass): Otherwise, perform a second query to the model. Prepend a special instruction indicating that the previous attempt showed uncertainty and encouraging more careful, step-by-step reasoning, then provide the same question again.
  4. 4.
    Final Answer: If a second pass was triggered, take the model’s second answer as the final answer for the question. Otherwise, use the first-pass answer.
Notably, in step 3 we do not give away any information about whether the first answer was correct, nor do we provide any external hints – we only indicate that the model itself was not confident and should rethink. This ensures we are not “cheating” by using ground truth; the model has to solve the problem on its own, albeit with a nudge to be more meticulous.
IV-C Reinforcement Prompt Design
A critical component of Reinforcement Inference is the design of the second-pass prompt. The goal is to make the model aware that its prior attempt was uncertain, while avoiding any task-specific hints or ground-truth leakage.
In our implementation, the second-pass input consists of three parts:
  1. 1.
    Original question block: the same question and options in the same structure as the first pass (e.g., Question: ..., Options: ..., Answer: ...).
  2. 2.
    Previous answer: a header line Your previous answer was: followed by the full first-round model output.
  3. 3.
    Instruction: a header line Instruction: followed by a fixed prompt indicating uncertainty (high entropy) and asking the model to discard the prior answer and recompute from first principles.
Importantly, this reprompting does not tell the model that it was wrong and does not provide the correct option. It only provides the previous output as context and asks for a fresh derivation under an explicit uncertainty cue.
IV-D Ensuring No Ground-Truth Leakage
A fundamental design principle for Reinforcement Inference is that the model should not accidentally get the correct answer through our intervention unless it can infer it by itself. We took several precautions to uphold this:
  • No answer revelation: We never show the model the correct answer, nor do we indicate whether its first answer was correct or incorrect.
  • Second pass includes the prior output, but no label: The second-pass input includes the model’s own first-pass raw output (“Your previous answer was: …”) to provide context. However, we never provide any signal about correctness (no “wrong”, no “correct”, and no ground-truth option).
  • Standardized instruction: The re-ask instruction text is fixed across all questions. It only indicates uncertainty (high entropy) and asks the model to discard the prior answer and recompute from first principles.
With these measures, the second-pass interaction is essentially the model solving the same question anew, but under a slight cognitive “pressure” to be careful. It’s analogous to giving a student the same problem again after telling them “you seemed unsure, try to be thorough now.” They still have to solve it themselves, but they know they should double-check their logic. Importantly, our evaluation treats the second attempt as the model’s final answer; we only count an answer as correct if the model actually produces the correct option on its own in that pass.
In summary, our method functions as a kind of self-reinforcement loop: the model’s own uncertainty triggers further reasoning by the model itself. No external hints or answers are given. This maintains the purity of the test scenario (important for fair benchmarking) and ensures that any improvement in accuracy comes from the model’s existing knowledge being better utilized, rather than from us covertly feeding it answers.
V Experimental Setup
V-A Model and Experiment Settings
We conducted experiments with DeepSeek-v3.2, a state-of-the-art LLM released in late 2025. DeepSeek-v3.2 is an open-source model with approximately 685 billion parameters, trained on a massive corpus of diverse multilingual and technical data [29]. It has demonstrated strong performance on a variety of benchmarks [29]. We chose DeepSeek-v3.2 for its strong reasoning ability, as we wanted a model that knows a lot but can still be challenged by difficult questions – an ideal candidate to test uncertainty-based reinforcement inference. Table I summarizes the inference parameters used across all experiments.
We used the model in a zero-shot setting with deterministic greedy decoding (temperature $t=0$) so that outputs and probabilities were stable across runs. The prompt format was kept consistent across first and second passes except for the added instruction in the second pass. Specifically, we formatted each multiple-choice question as a short prompt containing the question and the labeled options (A through J), and we instructed the model to output only the letter of the answer (e.g., “A”). This follows best practices from prior work to avoid variability in answer phrasing [31]. In the second pass, the same question format was used but with the reinforcement instruction prepended at the beginning of the prompt.
To extract the model’s answer probabilities, we used DeepSeek’s API to obtain log-probabilities for the output tokens. Since the answer in our setup is a single token (one of “A” through “J”), we could directly retrieve the probability the model assigned to each option at the answer position. These probabilities were normalized to sum to 1 over the ten options, and then we computed the entropy $H$ and Max $P$ from these values as described earlier.
TABLE I: Inference parameters.
Parameter Value
Model deepseek-chat
Temperature 0
Max tokens 4000
Top-pp 1
Frequency penalty 0
Presence penalty 0
Logprobs True
Top logprobs 20
V-B Dataset and Questions
We drew our evaluation questions from MMLU-Pro, which is an enhanced subset of the Massive Multitask Language Understanding benchmark focusing on professional or expert-level exams [32, 33]. MMLU is a collection of 57 subject areas (history, math, science, law, medicine, etc.) with questions ranging from elementary level to professional certification level [32]. The “Pro” subset specifically includes domains like medical board exams, law school exams, advanced mathematics, and other graduate-level or professional topics [35]. These questions are known to be especially challenging for LLMs – indeed, Hendrycks et al. (2021) noted that models often perform near chance on some professional subjects and, crucially, do not know when they are wrong [32, 36].
V-C Reference Results on MMLU-Pro
To contextualize the difficulty of MMLU-Pro, Table II reproduces a subset of the reference results reported in the MMLU-Pro paper (CoT, accuracies in %) [35].
TABLE II: A subset of reference MMLU-Pro results reported by [35] (accuracies in %).
Model Overall Math Physics Eng. Hist. Law Psych.
GPT-4o 72.6 76.1 74.7 55.0 70.1 51.0 79.2
Gemini-1.5-Pro 69.0 72.8 70.4 48.7 65.6 50.8 77.2
Claude-3-Opus 68.5 69.6 69.7 48.4 61.4 53.5 76.3
GPT-4-Turbo 63.7 62.8 61.0 35.9 67.7 51.2 78.3
Llama-3-70B-Instruct 56.2 54.0 49.6 43.6 56.9 39.9 70.2
From the MMLU-Pro repository, we evaluated 12,032 multiple-choice questions across 14 professional subjects: biology, business, chemistry, computer science, economics, engineering, health, history, law, math, other, philosophy, physics, and psychology [35]. Each question provided ten answer choices (A through J). All questions were drawn from the official test set of MMLU-Pro; the correct answers were hidden from the model (we only used them to check accuracy after the model produced its answer).
Out of these 12,032 questions, DeepSeek-v3.2 answered 7,306 correctly and 4,726 incorrectly on the first pass. This 60.72% accuracy is in line with expectations for a strong model on difficult zero-shot questions–leaving ample room for improvement. Using our entropy and max-probability thresholds, 7,347 questions (61.06%) were flagged as high uncertainty and received a second pass; the remaining 4,685 (38.94%) were accepted as confident and not re-asked.
V-D Procedure
Our empirical study consists of four experiments that share a common inference pipeline: for each question, we run a deterministic first pass to obtain (i) the model’s answer and (ii) token log-probabilities over the answer options, from which we compute the uncertainty metrics (entropy, MSP, and scoring-rule diagnostics). We then optionally run a second pass (re-ask) depending on the experimental condition.
Experiment I (effectiveness and causality). We evaluate our main method and two ablations on the same MMLU-Pro test set:
  • Baseline (one-shot): run only the first pass and use its answer as final.
  • Targeted re-asking (TR, main method): run the first pass, compute uncertainty, and trigger exactly one second pass when the trigger criterion is satisfied; the second-pass answer becomes the final answer.
  • Uniform re-asking (UR, ablation): run a second pass for every question (100%), regardless of uncertainty.
  • Prompt-only re-asking (ablation): run a single pass using the same “re-ask” prompt template as the second pass, but without performing an actual first-pass attempt (and thus without any uncertainty measurement or answer-conditioning). This isolates whether the observed gains could be explained purely by the stronger re-ask prompt wording itself.
Experiment II (generalization across model architectures and reasoning paradigms). We extend evaluation to additional state-of-the-art model families (DeepSeek-v3.2-Reasoner, Qwen3-30B-Instruct, and Qwen3-235B-Instruct) to test whether uncertainty-triggered re-asking transfers across (i) distinct reasoning paradigms and (ii) different model architectures and parameter scales, while maintaining inference parameter settings in Experiment I. We keep MMLU-Pro as the testbed for comparability and reuse the same static thresholds as in Experiment I ($\tau_{H}=1.3$ nats, $\tau_{\mathrm{MSP}}=0.4$) to probe robustness across models.
Experiment III (robustness / compute trade-off). We repeat the targeted re-asking policy over a grid of threshold pairs $(\tau_{H},\tau_{\mathrm{MSP}})$. For each pair, we record (i) the re-ask rate (fraction of questions triggering a second pass) and (ii) the resulting overall accuracy, producing an empirical accuracy–compute trade-off curve.
Experiment IV (cross-dataset transfer). We transfer the entropy threshold across tasks with different numbers of options by mapping a normalized entropy threshold from 10-way MMLU-Pro to the corresponding unnormalized threshold for 4-way MMLU. We then run the same targeted re-asking procedure on standard MMLU and report transfer performance.
Across all experiments, we evaluate final correctness using ground-truth labels only after the model has produced its outputs; no ground-truth information is ever shown to the model during inference.
V-E Statistical Tests
In addition to uncertainty metrics derived from the predictive distribution, we evaluate model confidence using proper scoring rules, which assess the quality of probabilistic predictions with respect to the ground-truth label. Proper scoring rules are designed such that the expected score is minimized when the predicted distribution matches the true data-generating distribution, thereby encouraging both accuracy and calibration.
We consider two widely used proper scoring rules: the Brier score and the negative log-likelihood (NLL). The Brier score measures the squared error between the predicted probability vector and the one-hot encoded ground-truth label, penalizing both overconfident incorrect predictions and underconfident correct predictions. The negative log-likelihood corresponds to the cross-entropy loss evaluated at the true label and grows unbounded as the predicted probability of the correct option approaches zero, making it particularly sensitive to extreme miscalibration. Together, these metrics provide complementary views of probabilistic quality: while entropy and maximum probability characterize intrinsic model uncertainty at inference time, Brier score and NLL quantify how well the model’s confidence aligns with empirical correctness.
[…]
$\chi^{2}=\frac{(|b-c|-1)^{2}}{b+c},$
which is approximately $\chi^{2}_{1}$-distributed under the null (continuity correction shown). We report the corresponding p-value for the high-uncertainty group.
Additionally, we report the raw improvement in accuracy (in percentage points) and compute an effect size for the change. Owing to the large evaluation set of 12,032 questions spanning diverse subjects, aggregate statistics naturally average out idiosyncratic fluctuations at the individual-item level, reducing the likelihood that the observed effects are driven by accidental alignment between a particular parameter setting and a small subset of questions. This scale enables robust paired tests and reliable subject-wise analysis.
VI Results
Experiments I evaluate effectiveness and causality (baseline vs. targeted re-asking vs. uniform re-asking). Experiments II evaluate generalization across model architectures and reasoning paradigms under fixed thresholds. Experiments III analyze robustness and the accuracy–compute trade-off via a 16-run threshold sweep. Experiments IV evaluate zero-shot transfer to standard MMLU by mapping the 10-way entropy threshold to a 4-way equivalent via normalized entropy.
VI-A Model Uncertainty vs. Answer Correctness (Baseline)
We first analyze the baseline one-shot pass and ask whether the model’s own uncertainty correlates with correctness. We find a clear separation: DeepSeek-v3.2 is typically more uncertain on questions it answers incorrectly. Figure 2 summarizes this baseline relationship between uncertainty (entropy and maximum probability) and correctness.
Figure 2: First-round distribution of entropy (nats) and MSP on the 12,032-question MMLU-Pro, illustrating the separation between correct and incorrect answers.
From our calculations on 12,032 questions (baseline pass), when the model answered incorrectly, its output distribution was typically much more spread out. The average entropy for correct answers was $0.844$ nats, compared to $1.567$ nats for incorrect answers. Similarly, the maximum probability assigned to the chosen answer was on average much higher for correct answers ($0.554$) than for incorrect answers ($0.249$). In short, baseline uncertainty is strongly predictive of mistakes.
TABLE III: Descriptive statistics of uncertainty metrics (MMLU-Pro, 12,032 questions). Values are mean $\pm$ standard deviation.
Round Metric Correct Incorrect
Baseline Entropy (nats) $0.844456\pm 0.763501$ $1.566893\pm 0.512095$
TR Entropy (nats) $0.940312\pm 0.881968$ $1.261358\pm 0.877025$
Baseline MSP $0.553804\pm 0.349872$ $0.249250\pm 0.206929$
TR MSP $0.539224\pm 0.365690$ $0.409469\pm 0.339603$
TABLE IV: Descriptive statistics of proper scoring rules (lower is better). Values are mean $\pm$ standard deviation.
Round Metric Correct Incorrect
Baseline Brier score $0.446196\pm 0.349872$ $0.965031\pm 0.322060$
TR Brier score $0.460776\pm 0.365690$ $1.214988\pm 0.486595$
Baseline NLL $0.844456\pm 0.763501$ $6.563630\pm 8.919788$
TR NLL $0.940312\pm 0.881968$ $11.663336\pm 10.542678$
TABLE V: Subject-wise re-asking coverage and accuracy on MMLU-Pro. “1st Acc.” and “2nd Acc.” are the overall accuracies within each subject (computed over all questions in that subject), using the first-pass answer vs. the final TR answer after uncertainty-triggered re-asking. “Re-asked” and “Re-asked (%)” report how often TR triggers within each subject; “Change” is the absolute improvement in percentage points.
Subject Total Re-asked Re-asked (%) 1st Acc. (overall, %) 2nd Acc. (overall, %) Change (pp)
Biology 717 599 83.54 65.13 92.47 +27.34
Business 789 199 25.22 80.61 87.33 +6.72
Chemistry 1132 505 44.61 78.98 88.16 +9.19
Computer Sci. 410 294 71.71 53.90 86.59 +32.68
Economics 844 715 84.72 52.25 87.44 +35.19
Engineering 969 570 58.82 75.44 80.08 +4.64
Health 818 346 42.30 76.41 80.56 +4.16
History 381 379 99.48 14.44 74.54 +60.10
Law 1101 1083 98.37 16.26 66.30 +50.05
Math 1351 738 54.63 57.29 92.75 +35.46
Other 924 308 33.33 80.19 80.30 +0.11
Philosophy 499 478 95.79 43.09 81.16 +38.08
Physics 1299 494 38.03 77.29 88.99 +11.70
Psychology 798 639 80.08 40.48 83.08 +42.61
In addition to entropy and MSP, proper scoring rules also separate correct from incorrect answers under the baseline pass: correct items have lower Brier score and lower NLL on average.
These baseline results support our key premise at scale: DeepSeek-v3.2’s hesitation signals—higher entropy and lower maximum probability—are strongly associated with mistakes on MMLU-Pro. This behavior is consistent with prior calibration findings on modern neural models [2, 6].
VI-B Impact of Reinforcement Inference on Accuracy (Targeted Re-asking, TR)
We now turn to the main evaluation: did the second-pass ”reinforcement” prompt actually lead to more correct answers on the flagged uncertain questions? The results confirm this hypothesis significantly.
Out of the 7,347 high-uncertainty questions where we triggered a second pass (TR), the model’s accuracy on that subset rose from 42.62% on first pass to 80.79% on second pass–an absolute gain of +38.18 percentage points. In terms of transitions: 2,912 questions improved from wrong to correct (39.64% of re-asked questions), 107 degraded from correct to wrong (1.46%), 3,024 stayed correct (41.16%), and 1,304 stayed wrong (17.75%). Thus the net change was +2,805 correct answers among the re-asked subset. The 4,685 questions that were not re-asked (those the model was confident about) maintained 89.11% accuracy on first pass and were left unchanged. Table V further breaks down re-asking coverage and before/after accuracy within each of the 14 subjects. The gains are observed across a wide range of subjects, supporting that TR is a cross-subject inference policy rather than a procedure tuned to fit any single subject.
[…]
$\chi^{2}=\frac{(|b-c|-1)^{2}}{b+c}=2604.3114,\qquad p<10^{-6}.$
In addition to statistical significance, the effect size is large: the re-asked subset improves by +38.18 percentage points (42.62% $\to$ 80.79%), with Cohen’s $h=0.8117$.
Across all 12,032 questions, overall accuracy rose from 60.72% on first pass to 84.03% after applying Reinforcement Inference, an absolute gain of +23.31 percentage points. The not re-asked subset (38.94% of questions) contributed its already-high 89.11% accuracy; the re-asked subset (61.06%) contributed the large improvement from 42.62% to 80.79%. A small number of initially correct answers (107) were flipped to wrong on second pass, but the net benefit was strongly positive.
Table IV reports the behavior of proper scoring rules (Brier score and negative log-likelihood, NLL) before and after uncertainty-triggered re-asking. While targeted re-asking (TR) substantially improves overall accuracy, we observe that for the remaining incorrect answers, both Brier score and NLL increase after TR. This effect is expected and reflects a change in the conditional distribution of errors rather than a degradation in model quality. Specifically, TR preferentially corrects errors associated with moderate uncertainty, leaving a residual set of incorrect instances that are intrinsically harder and for which the model assigns extremely low probability to the true option. As a result, the average loss under proper scoring rules—which strongly penalize near-zero probability assigned to the ground-truth label—increases for the incorrect subset after re-asking.
Overall, these results demonstrate that Reinforcement Inference succeeded in translating uncertainty awareness into concrete accuracy gains at scale. The model effectively self-corrected: when prompted about its own uncertainty, it was able to correctly solve a large fraction of the problems it initially missed. This suggests that the knowledge needed to answer many of those questions was latent in the model, and the second-pass prompt helped unlock it.
VI-C Ablation Study: Uniform Re-asking (UR)
To test whether the benefit of Reinforcement Inference comes from targeting uncertain questions or simply from re-asking more questions, we ran a uniform re-asking ablation (UR) in which we re-asked all 12,032 questions (100%) regardless of entropy or max probability. The results are summarized in Table VI.
TABLE VI: Targeted re-asking (TR) vs. uniform re-asking (UR).
Metric TR (Targeted) UR (Uniform)
Questions re-asked 7,347 (61.06%) 12,032 (100%)
Overall accuracy change +23.31% +23.63%
Final overall accuracy 84.03% 84.35%
[…]
TABLE VII: Summary of our inference-time evaluation on 12,032 MMLU-Pro items: mean entropy (nats) and mean maximum softmax probability (MaxProb) across inference rounds.
Round Entropy mean MaxProb mean
Baseline 1.1282 0.4342
TR 0.9916 0.5185
UR 1.1651 0.4463
When we re-asked every question (UR), overall accuracy improved by an additional +0.32 percentage points over the targeted approach (TR). In other words, re-asking only the 61.06% of questions flagged as uncertain achieved 98.6% of the maximum possible improvement (23.31% vs 23.63% gain). Re-asking the remaining 38.94% of questions (those where the model was already confident) provided negligible benefit. This confirms that entropy-based selection identifies exactly the questions that need re-asking: the model’s internal uncertainty is the key signal, not the volume of re-asking. It is not that the model cannot decide; it is that when the model answers under low confidence (high entropy), re-asking unlocks better reasoning. Addressing uncertainty is therefore central to addressing hallucination when the LLM is confused.
VI-D Ablation: Prompt-Only Re-asking vs. Uncertainty-Conditioned Re-asking
An alternative explanation for the observed gains is that they arise purely from prompt engineering effects, such as explicitly encouraging step-by-step reasoning or inducing a stronger chain-of-thought (CoT), rather than from the proposed uncertainty-conditioned re-asking mechanism. To disentangle these factors, we conduct an additional ablation that isolates the effect of the re-asking prompt itself.
Experimental setup. We compare the standard zero-shot baseline (DeepSeek-v3.2 evaluated on MMLU-Pro) against a prompt-only re-asking condition. In this ablation, the model is evaluated in a single pass under the same decoding and evaluation settings as the baseline, but with a modified system prompt:
The following are multiple choice questions (with answers) about {}. Your last output had high entropy, indicating uncertainty. Discard the prior answer and recompute the solution from first principles. Think step by step and then output the answer in the format of “The answer is (X)” at the end.
Critically, this condition does not include (i) the model’s first-round prediction, nor (ii) any actual uncertainty signal derived from a prior output, since no initial pass is performed. Thus, this ablation controls for both answer feedback and selective triggering, leaving only the effect of the re-asking instruction and CoT encouragement.
Results. The prompt-only re-asking condition underperforms the zero-shot baseline. That is, merely instructing the model to “discard the prior answer” and “recompute from first principles,” even with explicit step-by-step reasoning, does not improve—and in fact degrades—performance.
Implications. This result rules out the hypothesis that our improvements stem from the re-asking prompt or generic CoT elicitation alone. Instead, it provides evidence that the performance gains are driven by the interaction between (a) conditioning on the model’s own first-round answer and (b) selectively invoking re-asking based on a measured uncertainty signal. In the absence of a concrete prior attempt and a grounded justification for re-asking, the same prompt fails to yield benefits. This supports our central claim that uncertainty-aware, answer-conditioned re-asking, rather than prompt-level encouragement to “think harder,” is the key factor underlying the observed improvements.
VI-E Experiment II: Generalization Across Model Architectures and Reasoning Paradigms
Experimental setup. While our primary results demonstrate the efficacy of Reinforcement Inference on DeepSeek-v3.2, it remains unclear whether uncertainty-triggered self-correction is specific to one model family. We therefore evaluate three additional models with distinct characteristics: DeepSeek-v3.2-Reasoner (reasoning-oriented inference), Qwen3 30B, and Qwen3 235B (different architectures and scales). We keep MMLU-Pro as the testbed, while maintaining inference parameter settings in Experiment I, and reuse the static thresholds from Experiment I ($\tau_{H}=1.3$ nats, $\tau_{\mathrm{MSP}}=0.4$) to probe robustness.
Results. Table VIII reports overall (Ov.) and target-group (Tgt.; re-asked subset) statistics.
TABLE VIII: Generalization across model architectures on MMLU-Pro using static thresholds ($\tau_{H}=1.3$ nats, $\tau_{\mathrm{MSP}}=0.4$). This table summarizes the performance statistics for Qwen3 30B, Qwen3 235B, and DeepSeek V3.2 Reasoner. It reports the mean Entropy (Ent.) and mean Max Probability (Prob.) for the Overall dataset (Ov.) in Round 1 (R1) and Round 2 (R2). Additionally, it presents the Accuracy (Acc.) for both the Overall dataset and the Target Reasked group (Tgt.) across both rounds. The Reask Rate denotes the proportion of samples selected for re-evaluation.
Model Ov. R1 Ent. Ov. R1 Prob. Ov. R2 Ent. Ov. R2 Prob. Ov. R1 Acc. Ov. R2 Acc. Tgt. R1 Acc. Tgt. R2 Acc. Reask Rate
Qwen3 30B 1.2881 0.4292 1.3041 0.4300 0.7637 0.7725 0.7341 0.7494 0.5750
Qwen3 235B 1.4160 0.3753 1.4414 0.3752 0.8144 0.8258 0.7778 0.7953 0.6524
DeepSeek V3.2 0.2456 0.7546 0.1917 0.7872 0.7599 0.7870 0.3168 0.4461 0.2096
[…]
Implications. These cross-model differences suggest that static thresholds are not universally compute-efficient. A natural next step is model-specific calibration or dynamic uncertainty scaling (e.g., entropy normalization against a model’s baseline distribution), which can tune the accuracy–compute trade-off per architecture.
VI-F Experiment III: Threshold Sweep and Parameter Choices
To understand how sensitive Reinforcement Inference is to the uncertainty thresholds, we ran a sweep over 16 threshold pairs on MMLU-Pro:
$\tau_{H}\in\{0.7,0.9,1.1,1.3\}\ \text{nats},\qquad\tau_{\mathrm{MSP}}\in\{0.3,0.4,0.5,0.6\}.$
Figure 3: Accuracy–compute trade-off from the 16-run sweep. Each point is one threshold pair $(\tau_{H},\tau_{\mathrm{MSP}})$, plotted by its re-ask rate (x-axis) and final overall accuracy (y-axis).
How the Pareto curve is constructed. We treat each configuration as a point $(r,a)$, where $r$ is the re-ask rate (fraction of questions triggering a second pass; i.e., compute cost) and $a$ is the resulting final overall accuracy. A configuration $(r_{1},a_{1})$ is said to dominate $(r_{2},a_{2})$ if it achieves at least as much accuracy with no more compute ($r_{1}\leq r_{2}$ and $a_{1}\geq a_{2}$), and is strictly better in at least one dimension. The Pareto frontier in Figure 3 is the set of all non-dominated configurations. In practice, we obtain this frontier by sorting the 16 runs by increasing $r$ and scanning from low to high compute, keeping a point if its accuracy exceeds the best accuracy seen so far at any lower (or equal) re-ask rate.
The grid was chosen after a first-round trial analysis of the entropy/MSP distribution (Figure 2). Concretely, the values were selected to bracket the region where correct and incorrect instances overlap the most (mean entropy 0.845 vs. 1.567 nats; mean MSP 0.554 vs. 0.249), so that the sweep explores different trade-offs between compute (re-ask rate) and final accuracy.
Important caveat (threshold selection as an engineering aid). The 16-run sweep is guided by the first-round outcome distributions (i.e., after evaluation we can observe which instances are correct/incorrect and how entropy/MSP behave). The model itself never sees any ground-truth labels; however, using labeled outcomes on the same benchmark to motivate an operating region can be viewed as a form of “privileged” information for policy selection. We therefore treat this sweep primarily as an optimization/validation aid to identify stable, compute-efficient operating points and to inform future cross-domain and cross-dataset inference policies. A stricter protocol would select $\tau_{H},\tau_{\mathrm{MSP}}$ on a held-out validation split (or cross-validation) and only then report test performance.
Table IX summarizes the 16 runs (accuracy and re-ask rate). Unless otherwise noted, we use $\tau_{H}=0.8$ nats and $\tau_{\mathrm{MSP}}=0.6$ as our main configuration; this setting yields 84.03% final accuracy with 61.06% re-asking on MMLU-Pro.
TABLE IX: Threshold sweep on MMLU-Pro (16 runs). Each run uses trigger $\mathbb{I}[H>\tau_{H}\lor\mathrm{MSP}<\tau_{\mathrm{MSP}}]$.
Run $\tau_{H}$ (nats) $\tau_{\mathrm{MSP}}$ Re-ask (%) Acc. (%)
ent0.7_mp0.3 0.7 0.3 61.06 84.18
ent0.7_mp0.4 0.7 0.4 61.06 84.10
ent0.7_mp0.5 0.7 0.5 61.06 84.04
ent0.7_mp0.6 0.7 0.6 76.31 84.15
ent0.9_mp0.3 0.9 0.3 61.06 84.14
ent0.9_mp0.4 0.9 0.4 61.06 84.11
ent0.9_mp0.5 0.9 0.5 61.06 84.15
ent0.9_mp0.6 0.9 0.6 76.31 84.33
ent1.1_mp0.3 1.1 0.3 52.80 83.74
ent1.1_mp0.4 1.1 0.4 61.06 84.18
ent1.1_mp0.5 1.1 0.5 61.06 84.13
ent1.1_mp0.6 1.1 0.6 76.31 84.16
ent1.3_mp0.3 1.3 0.3 52.80 83.69
ent1.3_mp0.4 1.3 0.4 61.06 84.25
ent1.3_mp0.5 1.3 0.5 61.06 84.18
ent1.3_mp0.6 1.3 0.6 76.31 84.34
VI-G Experiment IV: Cross-Dataset Transfer: Zero-shot MMLU (4-way)
To test generalization beyond MMLU-Pro (10-way), we transferred the normalized entropy threshold implied by $\tau_{H}=1.3$ nats (10-way) to standard 4-way MMLU. Using
$\bar{H}_{\tau}=\frac{1.3}{\ln 10}\approx 0.565\quad\Rightarrow\quad\tau_{H}^{(K=4)}=\bar{H}_{\tau}\ln 4\approx 0.78\ \text{nats},$
with $\tau_{\mathrm{MSP}}=0.4$, we evaluated 14,042 zero-shot multiple-choice questions.
Overall, the baseline accuracy improved slightly from 91.134% to 91.184% (+0.050pp). Out of 57 subjects, 16 subjects improved, 11 subjects degraded, and the remainder were unchanged. This suggests that while uncertainty-triggered re-asking can yield large gains on the more difficult MMLU-Pro setting, its transfer to an already high-performing regime (standard MMLU at $\sim$91%) yields only minor net improvements and can introduce small regressions in some subjects.
TABLE X: Zero-shot MMLU transfer (4-way): normalized-entropy mapping from 10-way MMLU-Pro and MSP threshold kept fixed.
Metric Value
Dataset size 14,042 questions
$K$ (choices) 4
Entropy threshold $\tau_{H}$ (nats) 0.78
MSP threshold $\tau_{\mathrm{MSP}}$ 0.4
Baseline accuracy 91.134%
TR (Reinforcement Inference) accuracy 91.184%
Improved / Degraded subjects 16 / 11 (out of 57)
VII Discussion
Our findings showcase uncertainty-aware inference control in LLMs at scale. Reinforcement Inference allowed DeepSeek-v3.2 to emulate a human-like behavior: when unsure, pause and rethink. This method taps into the model’s internal uncertainty signals–which prior research suggests LLMs do possess to a degree [2, 6]–and leverages them in a practical way to improve performance without any additional training or external feedback. Below, we highlight several key takeaways about why the second-pass succeeds (e.g., recovering latent knowledge and changing decoding dynamics) and when it can fail (e.g., overconfident mistakes). These observations also help motivate the limitations and future directions discussed in the following sections.
  • Latent Knowledge Utilization: In many cases, the model did have the necessary knowledge or reasoning ability, as evidenced by the correct answer often having non-negligible probability even when the model answered incorrectly. The reinforcement prompt effectively said “you weren’t confident, so think harder.” This often caused the model to bring more thorough reasoning to bear, which uncovered the correct answer on the second try. For example, in the legal question mentioned earlier, the model clearly had some understanding of why option C was valid (it had assigned 25% to C initially), but only articulated that reasoning upon the second prompt. The first pass may have been a shallow pattern match or a guess based on surface cues, whereas the second pass induced a deeper analytical mode that let the model resolve its own uncertainty (akin to forcing a chain-of-thought when it matters).
  • Change of Decoding Dynamics: By alerting the model that it was uncertain and should reconsider, we implicitly instructed it to more strongly consider alternative options. When the model was unsure, it might have been nearly indifferent between two or three options. The second-pass prompt can break that tie by pushing the model to justify each option to itself, thereby tipping it toward the correct choice if the evidence supports it. In other words, the prompt influenced the decoding trajectory – instead of quickly settling on its top guess, the model “thinks out loud” a bit more. This is related to why chain-of-thought prompting helps [37, 38]: explaining or reasoning can clarify the model’s internal uncertainties. Our method induced a mini chain-of-thought only when needed, making it efficient.
[…]
  • Generality: While our experiment was on multiple-choice questions, the concept could extend to other task formats. In open-ended question answering, for instance, one could measure uncertainty by generating multiple different answers or by checking the probability of the answer the model gave. If uncertainty is high (e.g. the model’s top completion has low likelihood or there is high entropy over possible continuations), one could prompt the model to “explain and reconsider” before finalizing its answer. This would be analogous to our approach, though implementing it is trickier since open-ended answers don’t naturally produce a discrete probability distribution over options. Some recent work clusters semantically equivalent answers to define an entropy for open-generation settings [39, 40]. Our method might benefit those scenarios as well – for example, have the model first answer, then if it’s not confident (perhaps measured by the log-likelihood of that answer or by sampling variance), ask it to provide a justification and revise the answer. We suspect the same principle would hold: the model often has partial knowledge or clues that, when forced into the open by an explanation, can lead it to correct itself.
  • Efficiency Considerations: Reinforcement Inference adds some overhead–at most one extra model call for the flagged queries. In the targeted re-asking setting (TR), that was 7,347 extra calls for 12,032 queries (about a 61% increase in total queries run), achieving 99.6% of the accuracy gain of uniform re-asking (UR). This overhead is far lower than techniques like self-consistency, which might use 5–10 calls per query (a 500%–1000% overhead) to sample multiple reasoning paths for every question [9, 41]. It is also cheaper than strategies that always do chain-of-thought reasoning for every question. Our method is thus relatively lightweight. In interactive applications, this could translate to slightly longer latency only on the hard queries where the model was unsure, which might be a reasonable trade-off. One could even adjust the entropy threshold on the fly depending on desired speed vs. accuracy: for example, in a time-critical setting, only trigger the second pass if entropy is extremely high (i.e. only when the model is very unsure), whereas in a high-accuracy-required setting, use a lower threshold to double-check more often.
  • Comparison to “Step-Back” Prompting: Recently, some anecdotal reports and community forums have suggested manually prompting strategies where one asks an LLM to double-check an answer or take a “step back” when uncertain. Our approach formalizes and quantifies this idea by tying it to an actual uncertainty metric. Rather than always asking “Are you sure?” after every answer (which would be tedious and could even degrade performance if done when not needed), we selectively apply a step-back prompt when the model’s own probabilities indicate uncertainty. This selective trigger is a key novelty – it bridges confidence estimation with adaptive prompting. Essentially, we give the model a chance to reflect only when it hesitated in the first place. This focus makes our method both efficient and targeted, avoiding unnecessary second-guessing when the model was already confident (and likely correct).
VIII Limitations
Despite the encouraging results, our study has clear limitations. At the same time, they point to concrete next steps: strengthening uncertainty calibration across model families, extending entropy-triggered deliberation beyond multiple-choice tasks, and making the second-pass prompting policy more robust while remaining compute-aware. We discuss these limitations below.
  • Dependency on Model Calibration: A primary limitation of our approach is that its effectiveness is not uniform across architectures and appears to be deeply intertwined with a model’s training history, particularly concerning potential data contamination. We hypothesize that the minimal accuracy gains observed in the Qwen3 series, despite high re-asking rates, might indicate prior exposure to the MMLU-Pro dataset during its extensive post-training or alignment phases. In contrast, both the DeepSeek V3.2 and V3.2 Reasoner models demonstrated substantial performance improvements. This suggests these models likely had not been exposed to similar benchmark data during their training. Therefore, the observed accuracy gains can be more confidently attributed to the Reinforcement Inference process itself, which successfully guided the models to correct their reasoning in areas of genuine uncertainty. This positions our method as a tool for enhancing real-time reasoning rather than merely recovering memorized information.This highlights a critical challenge for generalization: the utility of our method is diminished on benchmarks that a model may have already been trained on. Its success is contingent on a model’s capacity for genuine generalization. Consequently, deploying this framework requires careful consideration of a model’s training provenance to ensure it can be meaningfully applied.
  • Multiple-Choice Focus: The current implementation is tailored to multiple-choice questions, where evaluating correctness is straightforward and the model’s output distribution is discrete over the given options. For open-ended tasks (free-form QA, code generation, etc.), applying a similar strategy requires more nuance in defining uncertainty and perhaps generating an alternative solution. It is doable – for example, one could sample multiple answers or use dropout to get an ensemble of outputs, and compute a form of entropy – but we did not explore that here. Thus, our results directly speak only to tasks with a closed set of answer options. Additional research is needed to generalize the approach to open-ended tasks.
  • Prompt Specificity: We used one particular phrasing for the second-pass prompt and observed good results, but we did not exhaustively compare alternatives. It’s possible that certain wordings could further improve effectiveness, or conversely that our prompt might not generalize to all contexts. For example, explicitly instructing the model “explain your reasoning before giving a final answer” in the second pass might yield even higher gains by ensuring it works through the problem (at the cost of longer outputs). We opted for a concise prompt to keep things simple. Future work could experiment with variations: there is a balance between nudging the model and over-constraining it, and we mostly gave it a light nudge. We also did not try adding few-shot examples of corrected reasoning; doing so might improve performance but would move away from our zero-shot paradigm.
  • Anchoring and prompt ablations: Our second-pass prompt includes the model’s own first-pass output (“Your previous answer was: …”) and explicitly states that the last output was uncertain (high entropy). While this still reveals no ground-truth label, it can introduce anchoring effects (the model may implicitly defend or over-correct relative to its prior output). We have included a prompt-only re-asking ablation (Section VI-D) showing that generic “think step-by-step” encouragement alone does not explain the gains; nevertheless, a more complete study should further ablate removing the prior output and/or removing the entropy cue to isolate each component.
  • No Adversarial or Trick Questions: Our evaluation did not include questions deliberately designed to fool the model’s uncertainty mechanism – for example, questions with a very appealing incorrect option that the model is 100% confident in (due to a common misconception), or cases where any hint of uncertainty leads the model to an incorrect line of reasoning. In such adversarial cases, our method would not trigger at all (because the model would be overconfident and not realize it’s wrong), or it might trigger but still fail to correct the error if the model’s reasoning is fundamentally flawed. We did not test on adversarial data, so our results reflect performance on normal challenging questions. Techniques to handle overconfident mistakes – perhaps by integrating an external verifier or by training the model to better recognize when it could be wrong – would be needed to address these scenarios. Essentially, Reinforcement Inference can only help when the model has some internal doubt or ambivalence; if the model is confidently wrong, other remedies are required (and identifying such cases is an open problem).
IX Future Work
The encouraging results from this initial study open up several directions for future work. In particular, we view Reinforcement Inference as an instance of a broader test-time control paradigm: using internal uncertainty estimates to allocate additional compute only when it is most likely to help. Realizing this vision robustly requires moving beyond a single benchmark and a single prompting template, and carefully characterizing when uncertainty signals are reliable versus misleading. The directions below therefore focus on generality, robustness, and system-level policies that make uncertainty-aware deliberation practical in real deployments:
  • Broader Evaluation: We plan to evaluate Reinforcement Inference on a broader range of tasks and models. For instance, testing with other advanced models like GPT-4 or newer open-source models would reveal whether they benefit similarly from uncertainty-triggered reasoning. This may require creative approaches if raw probabilities are not directly accessible (for some API models, one might prompt the model to estimate its confidence [11], although models can misreport confidence if not calibrated). Expanding to more benchmarks (e.g. other MMLU categories, reasoning-heavy datasets like GSM8K for math, or coding challenges) will help verify the generality of the method.
  • Dynamic Thresholds and Policies: In a deployed system, one could dynamically adjust the entropy threshold or use more sophisticated triggering policies. For example, a system could start with a certain $\tau$ but increase it if too many second passes are being triggered with little benefit (to save time), or lower it if the goal is maximum accuracy and extra compute is acceptable. One could even train a small classifier to predict when a second pass would likely be beneficial, based on features of the probability distribution and question. Automating the decision of when to invoke extra reasoning – essentially learning a meta-policy – is an interesting avenue.
  • Multi-Step Reinforcement: We limited our method to one extra pass. But what if the model is still uncertain after the second pass? One could consider iterating further: for example, a third pass with a prompt like “You are still unsure, try one more time considering each option carefully.” There may be diminishing returns, but a small number of iterations could potentially convert a few more hard cases. We would need to be cautious to not enter an endless loop; setting a maximum of perhaps two iterations beyond the initial attempt might be reasonable. Investigating how often a second pass remains uncertain (and if a third could help) is worthwhile future work.
  • Integrating Verification or Refusal: In critical applications, one might combine Reinforcement Inference with a final verification stage. For instance, if the model remains uncertain (high entropy) even after the reinforced reasoning attempt, perhaps it should then abstain or escalate to a human. In that sense, our approach could be part of a larger pipeline: (i) model gives an initial answer, (ii) if unsure, model tries a reinforced answer, (iii) if still unsure or the answer is suspect, the system refuses or hands off to a human. This would maximize the model’s chances to answer correctly on its own, but still maintain safety when the model truly doesn’t know the answer.
  • Applications in Generative Tasks: While we focused on Q&A, the idea of using uncertainty to trigger a self-correction could apply to generative tasks like story generation or code synthesis. For example, if a model is writing code and its internal probability distribution shows high entropy about the next steps (perhaps indicating confusion about how to proceed or a potential error), one could prompt it to “double-check the last steps” or generate an alternate solution. Similarly, in story or essay generation, if the model is uncertain, it might benefit from a prompt to rethink coherence or consistency. Exploring how to detect and act on uncertainty in creative generation is an intriguing direction.
  • Extending Reinforcement Inference to VLA and Robotics: An exciting direction for future work is to extend Reinforcement Inference beyond discrete answer selection to vision–language–action (VLA) and robotic control models [48, 47]. Modern VLA systems are increasingly decoder-based and autoregressive, generating action sequences token by token, which makes them a natural fit for uncertainty-triggered refinement.
    In this setting, entropy can be defined over the action distribution (e.g., discretized action tokens, spatial bins, or latent action representations), providing a measure of execution uncertainty. High spatial or action entropy corresponds to diffuse, unstable action proposals, while low entropy indicates confident, convergent trajectories. Reinforcement Inference could then be used to trigger additional refinement steps when the model remains uncertain, encouraging actions to progressively concentrate in high-certainty regions of the action space.
    This perspective aligns closely with human motor behavior: when performing a task such as grasping an object, initial movements are exploratory and uncertain, but as the hand approaches the target, actions naturally become more precise and less ambiguous. By explicitly conditioning additional action-generation passes on uncertainty signals, a VLA model may similarly improve reliability without relying on external rollouts, value functions, or explicit planning.
    Unlike traditional imitation learning, which optimizes for average behavior [50], or reinforcement learning, which depends on task-specific rewards, an entropy-constrained autoregressive approach focuses directly on reducing execution uncertainty. This idea is also compatible with modern visuomotor policy learning approaches (e.g., action diffusion) [49]. Exploring how such uncertainty-aware action refinement interacts with existing VLA and robotic learning paradigms is a promising direction for future research.
  • Entropy-Guided Training: Our work hints that models implicitly know when they’re unsure, but they are not optimized to act on that knowledge by default. A future direction is to incorporate uncertainty signals into the training process itself. For instance, one could add a auxiliary loss term that encourages the model to produce low entropy (high confidence) distributions only when it is correct, and high entropy when it is truly unsure. This would push the model to be confidently correct and appropriately hesitant when it doesn’t know, improving calibration. Recent research has stressed that LLMs must be taught to know what they don’t know [51]. Instead of only fine-tuning for better calibration post hoc [13, 28], one could train a model from the ground up with objectives that penalize incorrect high-confidence predictions. If successful, such a model would not only make fewer mistakes, but also be a perfect candidate for reinforcement inference techniques – its internal uncertainty signals would be even more trustworthy. In the long run, this line of research could lead to models that are jointly optimized for accuracy and for the correctness of their own confidence estimates.
X Conclusion
We introduced Reinforcement Inference, an inference-time approach that improves an LLM’s answers by leveraging its own uncertainty signals. Through experiments on 12,032 challenging multiple-choice questions from MMLU-Pro across 14 subjects, we demonstrated that when DeepSeek-v3.2 exhibited high predictive entropy—a proxy for hesitation—that uncertainty often signaled an incorrect first-pass answer, and that a second, more deliberate attempt could correct a large fraction of these errors without any retraining.
This work highlights several key insights. First, large language models like DeepSeek do have measurable internal confidence; in our tests at scale, wrong answers were usually accompanied by higher entropy and lower max probability than correct answers. This underscores the importance of developing techniques to read and use those internal signals for more reliable AI behavior. Second, a simple second-pass prompt (essentially telling the model, ”you were unsure, so be careful now”) can unlock better reasoning in the model. The success of this prompt shows how far prompt-based control can go in steering model cognition–in this case, nudging the model from a fast guess toward a more deliberative mode. Third, our method achieved substantial gains at scale: overall accuracy rose from 60.72% to 84.03% (+23.31%) while re-asking 61.06% of questions. A uniform re-asking ablation (UR) in which we re-asked 100% of questions yielded only +0.10% additional accuracy over targeted re-asking (TR), indicating that uncertainty-based triggering is the key–not the volume of re-asking–and that uncertainty is central to addressing hallucination when the LLM is confused. Nearly identical results with 39% less computation demonstrate both the efficiency and the correctness of our design. Crucially, all of this was done at inference time, preserving the model’s weights. This has practical implications: even when one cannot retrain a model (due to cost, or if using a fixed API model), it may still be possible to improve its reliability through intelligent post-processing or on-the-fly prompting of its outputs.
Reinforcement Inference contributes to the evolving toolbox for aligning and controlling LLM outputs [34]. It resonates with the idea of making models not just give an answer, but also reflect on how they arrived at it and whether they are satisfied with it. By operationalizing a form of self-reflection triggered by uncertainty, we move a step closer to LLMs that can know what they know–and know what they don’t. We hope this work inspires further research into uncertainty-aware AI and meta-cognitive prompting techniques, which together could make AI systems not only more accurate, but also more self-aware and trustworthy in their reasoning.
More broadly, our method is tightly coupled to the autoregressive nature of modern decoder-based models: because answers (and more generally, sequences) are generated token by token from a learned conditional distribution, entropy and related confidence measures arise naturally as first-class control signals during generation. Reinforcement Inference exploits this coupling by turning distributional uncertainty into an adaptive inference policy, allocating an additional pass only when the model’s own posterior indicates hesitation, thereby converting internal ambiguity into measurable and actionable compute. The same uncertainty-triggered refinement principle naturally extends beyond discrete answer selection to sequence-generation settings, including vision–language–action (VLA) models and robotic control. In such systems, spatial or action entropy can be interpreted as a proxy for execution uncertainty: high entropy corresponds to diffuse, unstable action proposals, whereas low entropy indicates confident, convergent trajectories. This perspective closely parallels human motor behavior, where early movements are exploratory and imprecise, progressively concentrating into precise actions as the hand approaches the target, motivating uncertainty-aware action refinement as a foundation for improving reliability in autoregressive VLA generation.
Stepping back, the broader message is that uncertainty is not merely a diagnostic; it is a control signal. Treating entropy as a first-class input to an adaptive inference policy provides a principled way to allocate compute, probe a model’s latent capability horizon, and improve reliability without changing model weights.

### v3

@@
 A well-known challenge is that models often lack calibrated confidence – they may not recognize when their answers are likely wrong [1, 2].
-This can lead to misleading outputs in high-stakes settings (for example, an LLM hallucinating a plausible-sounding but false explanation with full certainty [3, 4]).
+This can lead to misleading outputs in high-stakes settings (for example, an LLM hallucinating a plausible-sounding but false explanation with full certainty [3]).
 If an AI could know when it is unsure, it might mitigate errors by double-checking or deferring answers.
 Indeed, prior studies on tasks like multiple-choice QA have noted that an LLM’s softmax probabilities carry information about correctness [2, 4].
-Intuitively, a model that assigns one option nearly 100% probability is likely confident (and often correct), whereas if it spreads probability roughly uniformly across all options, it is essentially guessing [5, 2].
+Intuitively, a model that assigns one option nearly 100% probability is likely confident (and often correct), whereas if it spreads probability roughly uniformly across all options, it is essentially guessing [2].
 Recent evaluations confirm that wrong answers tend to have lower maximum softmax probability (MSP) than correct answers in well-trained models [2, 5].
@@
 Refusal might be safer in open-ended applications, but in many scenarios we do want the model to attempt an answer – just with more care if it was initially unsure.
-Humans do this naturally: if unsure of an answer, a person might pause and double-check their reasoning or try a different approach before responding.
+Humans do this naturally: if unsure of an answer, a person might pause and double-check their reasoning or try a different approach before responding [7].
 This inspires the question: Can an LLM similarly “think twice” when it hesitates, to produce a more accurate answer, all at inference time?
 Figure 1: System structure of Reinforcement Inference.
-In this paper, we propose an inference-time framework called Reinforcement Inference to realize this idea.
-The core concept is simple: the model’s own inferred uncertainty and MSP are used as a trigger for a second reasoning attempt.
-If the model’s first answer had high entropy – indicating confusion or low confidence – we prompt the model again with a special instruction that it was uncertain and should reason more thoroughly.
-Crucially, we do not reveal the correct answer or provide any external feedback; the model must rely on its existing knowledge, but is now explicitly guided to be careful.
-This procedure requires no additional training or parameter updates.
-It functions as a form of test-time control or meta-cognition for the model, akin to a student checking their work upon sensing uncertainty.
-As we will show, even this light-touch intervention can yield significant accuracy improvements on challenging questions that the model initially struggled with.
-Before diving into technical details, we emphasize the broader motivation of this work: modern LLM capability is often not fully realized by the default inference protocol.
-In many deployed settings—especially in professional and industrial workflows that require deterministic behavior—models are typically used with greedy decoding and a single forward pass.
-This “one-shot, greedy” regime can systematically under-estimate what a fixed model is actually capable of: many failures are not due to missing knowledge, but due to shallow first attempts, premature commitment, or unresolved internal ambiguity.
-Reinforcement Inference provides a practical way to probe and unlock a model’s latent capability horizon without any retraining.
-The large accuracy jump we observe is therefore not merely a reprompting trick; it is evidence that a substantial portion of the model’s reasoning capacity is available at inference time but remains untapped under the standard one-pass protocol.
-Methodologically, our key idea is to treat uncertainty as a first-class control signal during inference.
-Entropy and MSP expose when the model is effectively “guessing” under its own posterior, and we convert this signal into an adaptive compute policy: allocate a second reasoning pass only when the model itself indicates hesitation.
-This yields an efficiency–accuracy paradigm suited to real systems: predictable latency on easy cases, and deliberate reasoning only when needed.
-More broadly, we argue that entropy-aware inference suggests an entropy-aware training direction.
-For professional models where correctness and determinism matter, training objectives and decoding policies should explicitly constrain the relationship between correctness and confidence (e.g., penalizing high-confidence errors and encouraging calibrated decisiveness).
-Our work offers a concrete blueprint for such future designs: (i) use entropy/MSP to define when additional reasoning is warranted, (ii) quantify the resulting “capability gap” between one-pass greedy inference and uncertainty-conditioned deliberation, and (iii) use this gap as a diagnostic signal for both decoding and model optimization.
-In this sense, Reinforcement Inference is a step toward a broader paradigm in which uncertainty is not only measured, but also systematically acted upon.
+In this paper, we introduce Reinforcement Inference, an inference-time framework that lets an LLM use its own uncertainty to decide when to “think twice.” We run a deterministic first pass, monitor entropy and maximum softmax probability (MSP), and trigger a single additional, more deliberate attempt only when the model is uncertain.
+Crucially, this uses no labels or external feedback and requires no retraining.
+We argue that default one-shot greedy decoding can understate a fixed model’s true capability: in deterministic deployments, many failures reflect premature commitment under ambiguity rather than missing knowledge [7].
+Reinforcement Inference operationalizes this view by treating uncertainty as a first-class control signal, yielding an adaptive compute policy with predictable behavior on easy cases and extra deliberation only when needed.
+More broadly, our results motivate an entropy-aware training direction that explicitly shapes correctness–confidence alignment.
+The resulting “capability gap” between one-pass inference and uncertainty-conditioned deliberation can serve as a practical diagnostic for system design and model optimization.
 II Contributions
@@
 III-A Confidence and Uncertainty in LLMs
-There is growing interest in measuring how well an LLM’s predicted probabilities reflect its actual likelihood of being correct [2, 15], including recent work on whether LLMs can reason explicitly about their own confidence [44].
-Maximum softmax probability (MSP) has been used as a simple proxy for model confidence, with studies showing that lower MSP correlates with wrong answers and even with out-of-distribution inputs in both vision and NLP models [2, 15].
-For instance, Plaut et al.
-(2024) provide strong evidence that in multiple-choice QA, an LLM’s softmax distribution can predict whether its answer is correct [16, 17].
-They leveraged this by allowing the model to abstain on low-confidence answers, improving overall accuracy by avoiding some mistakes [46, 18, 17].
-Other works explore more complex uncertainty quantification methods (Bayesian approaches, ensembles, or probing internal states [17, 19]) and calibrating models via fine-tuning to output better confidence estimates [20, 21].
-Our work aligns with these efforts in that we use the model’s own probability distribution as an uncertainty signal.
-However, instead of training the model to be calibrated or to abstain, we use uncertainty to trigger a self-correction process at inference time.
+Recent work studies how well an LLM’s predictive probabilities reflect correctness and whether models can explicitly reason about their own confidence.
+Maximum softmax probability (MSP) is a common proxy and tends to be lower for incorrect and out-of-distribution inputs [1].
+Studies show that softmax distributions can predict correctness in multiple-choice QA and can be used for selective abstention [2, 6].
+Related work studies Bayesian/ensemble uncertainty estimation and calibration via fine-tuning [5, 8].
+We likewise use the model’s output distribution as an uncertainty signal, but rather than training calibration or abstention, we trigger self-correction at inference time.
 III-B Chain-of-Thought and Self-Consistency
-Prompting strategies like chain-of-thought (CoT) encourage models to generate intermediate reasoning steps, which often boosts performance on complex tasks [8].
-Building on CoT, the self-consistency approach samples multiple independent reasoning paths and then selects the most common answer among them [8, 9].
-Wang et al.
-(2023) showed that self-consistency significantly improves accuracy on math and commonsense questions by effectively ensembling reasoning outcomes [9].
-While powerful, self-consistency incurs substantial computational cost (many model forward passes per query) and always uses multiple attempts even if the model was confident on the first try.
-In contrast, our method performs a second pass selectively – only for cases where the first pass was uncertain.
-Thus, Reinforcement Inference can be seen as a targeted, single-step variant of reasoning augmentation.
-It resembles asking the model to “try a different line of reasoning” but only when warranted by low confidence, saving computation on easier queries.
+Chain-of-thought (CoT) prompting can improve performance by eliciting intermediate reasoning [9].
+Self-consistency extends CoT by sampling multiple reasoning paths and selecting the majority answer [10], yielding sizable gains on tasks such as math and commonsense reasoning.
+However, it is compute-intensive and uses multiple attempts regardless of confidence.
+In contrast, Reinforcement Inference triggers only a single additional pass and does so selectively when the first pass is uncertain, saving compute on easy queries.
 III-C Verifier Models and Answer Re-Ranking
-Another line of work improves answer correctness by introducing a verification stage.
-One paradigm is to train a separate verifier classifier that judges whether the model’s answer is correct (sometimes using explanations or retrieved evidence), and then use that to accept or reject answers [10, 11].
-Other approaches prompt the model itself to check its answer.
-For example, Chain-of-Verification (CoVe) has the model generate an initial answer, then pose follow-up verification questions to itself, answer those, and finally produce a revised answer [22, 23].
-This multi-step deliberation can reduce factual hallucinations [24, 25].
-Similarly, self-critique techniques have the model reflect on and potentially correct its first answer by analyzing its own output for errors [26, 27].
-These verification-based methods improve reliability, but they often rely on complex prompt programs or fine-tuned models for critique.
-Our approach is much simpler: we do not generate explicit verification questions or use a separate model; we just give a single additional prompt reminding the model to be careful.
-In effect, Reinforcement Inference is a lightweight self-reflection step triggered by uncertainty, rather than a full-fledged self-dialogue or learned verifier.
+Another line improves correctness via explicit verification.
+Some methods train separate verifiers to accept/reject model answers, while others prompt the model to verify itself: Chain-of-Verification (CoVe) iteratively generates an answer, asks and answers follow-up checks, and revises the output, reducing hallucinations [11].
+Related self-critique approaches similarly analyze and refine the initial output.
+Compared to these multi-step pipelines, Reinforcement Inference is lightweight: a single uncertainty-triggered second pass without learned verifiers or explicit verification questions.
 III-D Supervised Retraining for Reasoning
-A straightforward way to improve a model’s accuracy or calibration is additional training.
-Fine-tuning on high-quality reasoning data (e.g.
-with CoT exemplars or human explanations) can make a model more reliable on tricky problems.
-Reinforcement Learning from Human Feedback (RLHF) is also widely used to align models to preferred answers and avoid obvious mistakes.
-For example, recent work has fine-tuned LLMs specifically to output better calibrated uncertainty estimates or to know when to say “I don’t know” [13, 28].
-However, these methods require collecting labeled data or rewards and then retraining the model – which may not be feasible for very large or closed-source models.
-In contrast, our method introduces no new training at all: it works as an inference-time plug-in that could be applied to any pre-trained model (even black-box API models, provided one can extract a confidence measure like probabilities or logits).
+A common way to improve accuracy or calibration is additional training or fine-tuning on reasoning data or aligning with RLHF.
+Recent work has also fine-tuned LLMs to output better-calibrated uncertainty or to abstain (“I don’t know”) [12].
+However, such approaches require labeled data/rewards and retraining, which can be impractical for very large or closed-source models.
+In contrast, our method requires no training: it is an inference-time plug-in applicable to any model where a confidence signal (e.g., probabilities/logits) can be obtained.
 IV Methodology
@@
 From a Bayesian decision perspective, if the model’s posterior over options is $P$, the Bayes-optimal 0–1 decision is $\hat{y}=\arg\max_{i}P_{i}$, and $1-\mathrm{MSP}(P)$ upper-bounds the posterior risk of choosing $\hat{y}$ under the model’s own beliefs.
-Motivated by calibration theory, we also define two proper scoring rules [45], which we report as secondary diagnostics in Section VI-A:
+Inspired by classical post-hoc calibration methods such as Platt scaling [13], we also define two proper scoring rules [14], which we report as secondary diagnostics in Section VI-A:
   • Brier score: for one-hot label $y$,
@@
 To test this hypothesis, we ran DeepSeek-v3.2 on a large set of multiple-choice questions and extracted the model’s answer distribution over the provided options from token log-probabilities (converted to normalized probabilities).
-We then computed $H$ and Max $P$ for each question and compared these uncertainty metrics between items that were answered correctly vs.
+We then computed $H$ and $\mathrm{MSP}$ for each question and compared these uncertainty metrics between items that were answered correctly vs.
 incorrectly post hoc.
@@
 This ensures we are not “cheating” by using ground truth; the model has to solve the problem on its own, albeit with a nudge to be more meticulous.
-IV-C Reinforcement Prompt Design
+IV-C Second-Pass Prompting Protocol
 A critical component of Reinforcement Inference is the design of the second-pass prompt.
-The goal is to make the model aware that its prior attempt was uncertain, while avoiding any task-specific hints or ground-truth leakage.
+The goal is to make the model aware that its prior attempt was uncertain, while avoiding any task-specific hints.
 In our implementation, the second-pass input consists of three parts:
   1. 1.
-    Original question block: the same question and options in the same structure as the first pass (e.g., Question: ..., Options: ..., Answer: ...).
+    Original question block: the same question and options in the same structure as the first pass (e.g., Question: ..., Options: ...).
   2. 2.
@@
     Instruction: a header line Instruction: followed by a fixed prompt indicating uncertainty (high entropy) and asking the model to discard the prior answer and recompute from first principles.
-Importantly, this reprompting does not tell the model that it was wrong and does not provide the correct option.
-It only provides the previous output as context and asks for a fresh derivation under an explicit uncertainty cue.
-IV-D Ensuring No Ground-Truth Leakage
-A fundamental design principle for Reinforcement Inference is that the model should not accidentally get the correct answer through our intervention unless it can infer it by itself.
-We took several precautions to uphold this:
-  • No answer revelation: We never show the model the correct answer, nor do we indicate whether its first answer was correct or incorrect.
-  • Second pass includes the prior output, but no label: The second-pass input includes the model’s own first-pass raw output (“Your previous answer was: …”) to provide context.
-However, we never provide any signal about correctness (no “wrong”, no “correct”, and no ground-truth option).
-  • Standardized instruction: The re-ask instruction text is fixed across all questions.
-It only indicates uncertainty (high entropy) and asks the model to discard the prior answer and recompute from first principles.
-With these measures, the second-pass interaction is essentially the model solving the same question anew, but under a slight cognitive “pressure” to be careful.
-It’s analogous to giving a student the same problem again after telling them “you seemed unsure, try to be thorough now.” They still have to solve it themselves, but they know they should double-check their logic.
-Importantly, our evaluation treats the second attempt as the model’s final answer; we only count an answer as correct if the model actually produces the correct option on its own in that pass.
-In summary, our method functions as a kind of self-reinforcement loop: the model’s own uncertainty triggers further reasoning by the model itself.
-No external hints or answers are given.
-This maintains the purity of the test scenario (important for fair benchmarking) and ensures that any improvement in accuracy comes from the model’s existing knowledge being better utilized, rather than from us covertly feeding it answers.
+Importantly, the reprompt never tells the model it was wrong and never reveals the correct option.
+We only include the first-pass output as context and request a fresh derivation under an explicit uncertainty cue, so the second pass re-solves the same question with added pressure to be careful (analogous to asking a student to redo a problem after noting hesitation).
+In summary, this is a self-reinforcement loop: uncertainty triggers additional reasoning without external hints, preserving a clean evaluation and attributing gains to better use of the model’s existing knowledge rather than information leakage.
 V Experimental Setup
@@
 We conducted experiments with DeepSeek-v3.2, a state-of-the-art LLM released in late 2025.
-DeepSeek-v3.2 is an open-source model with approximately 685 billion parameters, trained on a massive corpus of diverse multilingual and technical data [29].
-It has demonstrated strong performance on a variety of benchmarks [29].
-We chose DeepSeek-v3.2 for its strong reasoning ability, as we wanted a model that knows a lot but can still be challenged by difficult questions – an ideal candidate to test uncertainty-based reinforcement inference.
+It is an open-source model with approximately 685 billion parameters, trained on a massive multilingual and technical corpus and strong across benchmarks [15].
+We chose it as a strong yet challenging testbed for uncertainty-based Reinforcement Inference.
 Table I summarizes the inference parameters used across all experiments.
-We used the model in a zero-shot setting with deterministic greedy decoding (temperature $t=0$) so that outputs and probabilities were stable across runs.
-The prompt format was kept consistent across first and second passes except for the added instruction in the second pass.
-Specifically, we formatted each multiple-choice question as a short prompt containing the question and the labeled options (A through J), and we instructed the model to output only the letter of the answer (e.g., “A”).
-This follows best practices from prior work to avoid variability in answer phrasing [31].
-In the second pass, the same question format was used but with the reinforcement instruction prepended at the beginning of the prompt.
-To extract the model’s answer probabilities, we used DeepSeek’s API to obtain log-probabilities for the output tokens.
-Since the answer in our setup is a single token (one of “A” through “J”), we could directly retrieve the probability the model assigned to each option at the answer position.
-These probabilities were normalized to sum to 1 over the ten options, and then we computed the entropy $H$ and Max $P$ from these values as described earlier.
+We evaluate in a zero-shot, deterministic setting with greedy decoding (temperature $t=0$) to ensure stable outputs and probabilities.
+Both passes use the same multiple-choice format (question + options A–J) and prompt the model to output answering process and option letter [16]; the second pass simply prepends the reinforcement instruction.
+To obtain answer probabilities, we query the API for token-level log-probabilities.
+Because the answer is a single token (A–J), we directly read out the probability of each option, normalize over the ten choices, and compute the entropy $H$ and $MSP$ as described earlier.
 TABLE I: Inference parameters.
 Parameter Value
-Model deepseek-chat
+Model deepseek-v3.2
 Temperature 0
@@
 We drew our evaluation questions from MMLU-Pro, which is an enhanced subset of the Massive Multitask Language Understanding benchmark focusing on professional or expert-level exams [17, 18].
-MMLU is a collection of 57 subject areas (history, math, science, law, medicine, etc.) with questions ranging from elementary level to professional certification level [32].
-The “Pro” subset specifically includes domains like medical board exams, law school exams, advanced mathematics, and other graduate-level or professional topics [35].
-These questions are known to be especially challenging for LLMs – indeed, Hendrycks et al.
-(2021) noted that models often perform near chance on some professional subjects and, crucially, do not know when they are wrong [32, 36].
+MMLU is a collection of 57 subject areas (history, math, science, law, medicine, etc.) with questions ranging from elementary level to professional certification level.
+The “Pro” subset specifically includes domains like medical board exams, law school exams, advanced mathematics, and other graduate-level or professional topics.
+These questions are known to be especially challenging for LLMs – models often perform near chance on some professional subjects and, crucially, do not know when they are wrong [19].
 V-C Reference Results on MMLU-Pro
@@
 V-E Statistical Tests
-In addition to uncertainty metrics derived from the predictive distribution, we evaluate model confidence using proper scoring rules, which assess the quality of probabilistic predictions with respect to the ground-truth label.
-Proper scoring rules are designed such that the expected score is minimized when the predicted distribution matches the true data-generating distribution, thereby encouraging both accuracy and calibration.
-We consider two widely used proper scoring rules: the Brier score and the negative log-likelihood (NLL).
-The Brier score measures the squared error between the predicted probability vector and the one-hot encoded ground-truth label, penalizing both overconfident incorrect predictions and underconfident correct predictions.
-The negative log-likelihood corresponds to the cross-entropy loss evaluated at the true label and grows unbounded as the predicted probability of the correct option approaches zero, making it particularly sensitive to extreme miscalibration.
-Together, these metrics provide complementary views of probabilistic quality: while entropy and maximum probability characterize intrinsic model uncertainty at inference time, Brier score and NLL quantify how well the model’s confidence aligns with empirical correctness.
+In addition to uncertainty metrics from the predictive distribution, we assess confidence using proper scoring rules, which evaluate probabilistic predictions against the ground-truth label and are minimized in expectation by the true distribution.
+We use the Brier score and negative log-likelihood (NLL).
+The Brier score is the squared error between the predicted probability vector and the one-hot label, while NLL (cross-entropy at the true label) strongly penalizes assigning low probability to the correct option.
+These complement entropy/MSP by measuring how well confidence aligns with empirical correctness.
 […]
@@
 We find a clear separation: DeepSeek-v3.2 is typically more uncertain on questions it answers incorrectly.
-Figure 2 summarizes this baseline relationship between uncertainty (entropy and maximum probability) and correctness.
+Figure 2 summarizes this baseline relationship between uncertainty (entropy and maximum probability) and correctness [15].
 Figure 2: First-round distribution of entropy (nats) and MSP on the 12,032-question MMLU-Pro, illustrating the separation between correct and incorrect answers.
@@
 A small number of initially correct answers (107) were flipped to wrong on second pass, but the net benefit was strongly positive.
-Table IV reports the behavior of proper scoring rules (Brier score and negative log-likelihood, NLL) before and after uncertainty-triggered re-asking.
-While targeted re-asking (TR) substantially improves overall accuracy, we observe that for the remaining incorrect answers, both Brier score and NLL increase after TR.
-This effect is expected and reflects a change in the conditional distribution of errors rather than a degradation in model quality.
-Specifically, TR preferentially corrects errors associated with moderate uncertainty, leaving a residual set of incorrect instances that are intrinsically harder and for which the model assigns extremely low probability to the true option.
-As a result, the average loss under proper scoring rules—which strongly penalize near-zero probability assigned to the ground-truth label—increases for the incorrect subset after re-asking.
-Overall, these results demonstrate that Reinforcement Inference succeeded in translating uncertainty awareness into concrete accuracy gains at scale.
-The model effectively self-corrected: when prompted about its own uncertainty, it was able to correctly solve a large fraction of the problems it initially missed.
-This suggests that the knowledge needed to answer many of those questions was latent in the model, and the second-pass prompt helped unlock it.
+Table IV reports proper scoring rules (Brier score and negative log-likelihood, NLL) before and after uncertainty-triggered re-asking.
+Although targeted re-asking (TR) improves accuracy, Brier and NLL can increase on the remaining incorrect cases.
+This is expected: TR mainly fixes moderate-uncertainty mistakes, leaving a smaller set of intrinsically hard errors where the model assigns near-zero probability to the true option, which these losses penalize strongly.
+Overall, Reinforcement Inference translates uncertainty awareness into large accuracy gains via effective self-correction, suggesting that many initially missed answers are latent and can be recovered with a second-pass prompt.
 VI-C Ablation Study: Uniform Re-asking (UR)
@@
 Think step by step and then output the answer in the format of “The answer is (X)” at the end.
-Critically, this condition does not include (i) the model’s first-round prediction, nor (ii) any actual uncertainty signal derived from a prior output, since no initial pass is performed.
+This condition does not include (i) the model’s first-round prediction, nor (ii) any actual uncertainty signal derived from a prior output, since no initial pass is performed.
 Thus, this ablation controls for both answer feedback and selective triggering, leaving only the effect of the re-asking instruction and CoT encouragement.
@@
 Implications.
-This result rules out the hypothesis that our improvements stem from the re-asking prompt or generic CoT elicitation alone.
-Instead, it provides evidence that the performance gains are driven by the interaction between (a) conditioning on the model’s own first-round answer and (b) selectively invoking re-asking based on a measured uncertainty signal.
-In the absence of a concrete prior attempt and a grounded justification for re-asking, the same prompt fails to yield benefits.
-This supports our central claim that uncertainty-aware, answer-conditioned re-asking, rather than prompt-level encouragement to “think harder,” is the key factor underlying the observed improvements.
+This indicates that gains do not come from the prompt or generic CoT elicitation alone; they require uncertainty-triggered, selective re-asking conditioned on a concrete first-round attempt.
+In short, the key is uncertainty-aware, answer-conditioned re-asking rather than urging the model to “think harder.”
 VI-E Experiment II: Generalization Across Model Architectures and Reasoning Paradigms
@@
 TABLE VIII: Generalization across model architectures on MMLU-Pro using static thresholds ($\tau_{H}=1.3$ nats, $\tau_{\mathrm{MSP}}=0.4$).
-This table summarizes the performance statistics for Qwen3 30B, Qwen3 235B, and DeepSeek V3.2 Reasoner.
+This table summarizes the performance statistics for Qwen3 30B, Qwen3 235B, and DeepSeek-v3.2-Reasoner.
 It reports the mean Entropy (Ent.) and mean Max Probability (Prob.) for the Overall dataset (Ov.) in Round 1 (R1) and Round 2 (R2).
@@
 Model Ov. R1 Ent. Ov. R1 Prob. Ov. R2 Ent. Ov. R2 Prob. Ov. R1 Acc. Ov. R2 Acc. Tgt. R1 Acc. Tgt. R2 Acc. Reask Rate
-Qwen3 30B 1.2881 0.4292 1.3041 0.4300 0.7637 0.7725 0.7341 0.7494 0.5750
-Qwen3 235B 1.4160 0.3753 1.4414 0.3752 0.8144 0.8258 0.7778 0.7953 0.6524
-DeepSeek V3.2 0.2456 0.7546 0.1917 0.7872 0.7599 0.7870 0.3168 0.4461 0.2096
+Qwen3 30B 1.2881 0.4292 1.3041 0.4300 76.37% 77.25% 73.41% 74.94% 57.50%
+Qwen3 235B 1.4160 0.3753 1.4414 0.3752 81.44% 82.58% 77.78% 79.53% 65.24%
+DeepSeek V3.2 0.2456 0.7546 0.1917 0.7872 75.99% 78.70% 31.68% 44.61% 20.96%
 […]
@@
 How the Pareto curve is constructed.
-We treat each configuration as a point $(r,a)$, where $r$ is the re-ask rate (fraction of questions triggering a second pass; i.e., compute cost) and $a$ is the resulting final overall accuracy.
-A configuration $(r_{1},a_{1})$ is said to dominate $(r_{2},a_{2})$ if it achieves at least as much accuracy with no more compute ($r_{1}\leq r_{2}$ and $a_{1}\geq a_{2}$), and is strictly better in at least one dimension.
-The Pareto frontier in Figure 3 is the set of all non-dominated configurations.
-In practice, we obtain this frontier by sorting the 16 runs by increasing $r$ and scanning from low to high compute, keeping a point if its accuracy exceeds the best accuracy seen so far at any lower (or equal) re-ask rate.
-The grid was chosen after a first-round trial analysis of the entropy/MSP distribution (Figure 2).
-Concretely, the values were selected to bracket the region where correct and incorrect instances overlap the most (mean entropy 0.845 vs.
-1.567 nats; mean MSP 0.554 vs.
-0.249), so that the sweep explores different trade-offs between compute (re-ask rate) and final accuracy.
+Each threshold pair is a point $(r,a)$, where $r$ is the re-ask rate (compute) and $a$ is final accuracy.
+We define dominance as $r_{1}\leq r_{2}$ and $a_{1}\geq a_{2}$ (strict in at least one), and the Pareto frontier in Figure 3 is the set of non-dominated points.
+We compute it by sorting runs by increasing $r$ and retaining a point whenever it achieves a new best accuracy at that (or lower) compute.
+The 16-point grid is guided by the first-round entropy/MSP distributions (Figure 2) to bracket where correct and incorrect cases overlap most, enabling a focused sweep over the compute–accuracy trade-off.
 Important caveat (threshold selection as an engineering aid).
@@
 with $\tau_{\mathrm{MSP}}=0.4$, we evaluated 14,042 zero-shot multiple-choice questions.
-Overall, the baseline accuracy improved slightly from 91.134% to 91.184% (+0.050pp).
-Out of 57 subjects, 16 subjects improved, 11 subjects degraded, and the remainder were unchanged.
-This suggests that while uncertainty-triggered re-asking can yield large gains on the more difficult MMLU-Pro setting, its transfer to an already high-performing regime (standard MMLU at $\sim$91%) yields only minor net improvements and can introduce small regressions in some subjects.
+Overall, accuracy changes only marginally on standard MMLU: 91.134% $\to$ 91.184% (+0.050pp).
+Across 57 subjects, 16 improve and 11 degrade, with the rest unchanged.
+Thus, while uncertainty-triggered re-asking yields large gains on harder MMLU-Pro, its transfer to an already strong regime ($\sim$91%) provides only minor net benefit and may cause small subject-level regressions.
 TABLE X: Zero-shot MMLU transfer (4-way): normalized-entropy mapping from 10-way MMLU-Pro and MSP threshold kept fixed.
@@
 VII Discussion
-Our findings showcase uncertainty-aware inference control in LLMs at scale.
-Reinforcement Inference allowed DeepSeek-v3.2 to emulate a human-like behavior: when unsure, pause and rethink.
-This method taps into the model’s internal uncertainty signals–which prior research suggests LLMs do possess to a degree [2, 6]–and leverages them in a practical way to improve performance without any additional training or external feedback.
-Below, we highlight several key takeaways about why the second-pass succeeds (e.g., recovering latent knowledge and changing decoding dynamics) and when it can fail (e.g., overconfident mistakes).
-These observations also help motivate the limitations and future directions discussed in the following sections.
-  • Latent Knowledge Utilization: In many cases, the model did have the necessary knowledge or reasoning ability, as evidenced by the correct answer often having non-negligible probability even when the model answered incorrectly.
-The reinforcement prompt effectively said “you weren’t confident, so think harder.” This often caused the model to bring more thorough reasoning to bear, which uncovered the correct answer on the second try.
-For example, in the legal question mentioned earlier, the model clearly had some understanding of why option C was valid (it had assigned 25% to C initially), but only articulated that reasoning upon the second prompt.
-The first pass may have been a shallow pattern match or a guess based on surface cues, whereas the second pass induced a deeper analytical mode that let the model resolve its own uncertainty (akin to forcing a chain-of-thought when it matters).
-  • Change of Decoding Dynamics: By alerting the model that it was uncertain and should reconsider, we implicitly instructed it to more strongly consider alternative options.
-When the model was unsure, it might have been nearly indifferent between two or three options.
-The second-pass prompt can break that tie by pushing the model to justify each option to itself, thereby tipping it toward the correct choice if the evidence supports it.
-In other words, the prompt influenced the decoding trajectory – instead of quickly settling on its top guess, the model “thinks out loud” a bit more.
-This is related to why chain-of-thought prompting helps [37, 38]: explaining or reasoning can clarify the model’s internal uncertainties.
-Our method induced a mini chain-of-thought only when needed, making it efficient.
+Our findings demonstrate uncertainty-aware inference control at scale.
+Reinforcement Inference enables DeepSeek-v3.2 to behave more like a careful human: when uncertain, pause and rethink.
+It leverages internal uncertainty signals [2, 5] to improve performance without retraining or external feedback.
+Below we summarize why the second pass helps and when it can fail, motivating the limitations and future directions that follow.
+  • Latent Knowledge Utilization: The correct option often has non-negligible probability even when the first-pass answer is wrong, suggesting latent knowledge is present but underused.
+The reinforcement cue (“you weren’t confident, so think harder”) often elicits deeper reasoning and recovers the correct answer.
+  • Change of Decoding Dynamics: When the model is nearly indifferent among a few options, the second-pass cue encourages considering alternatives and justifying choices, often breaking ties in favor of the correct option.
+This echoes why CoT can help [20, 21]: explicit reasoning can resolve internal ambiguity.
+RI induces this extra deliberation only on uncertain cases.
 […]
-  • Generality: While our experiment was on multiple-choice questions, the concept could extend to other task formats.
-In open-ended question answering, for instance, one could measure uncertainty by generating multiple different answers or by checking the probability of the answer the model gave.
-If uncertainty is high (e.g.
-the model’s top completion has low likelihood or there is high entropy over possible continuations), one could prompt the model to “explain and reconsider” before finalizing its answer.
-This would be analogous to our approach, though implementing it is trickier since open-ended answers don’t naturally produce a discrete probability distribution over options.
-Some recent work clusters semantically equivalent answers to define an entropy for open-generation settings [39, 40].
-Our method might benefit those scenarios as well – for example, have the model first answer, then if it’s not confident (perhaps measured by the log-likelihood of that answer or by sampling variance), ask it to provide a justification and revise the answer.
-We suspect the same principle would hold: the model often has partial knowledge or clues that, when forced into the open by an explanation, can lead it to correct itself.
-  • Efficiency Considerations: Reinforcement Inference adds some overhead–at most one extra model call for the flagged queries.
-In the targeted re-asking setting (TR), that was 7,347 extra calls for 12,032 queries (about a 61% increase in total queries run), achieving 99.6% of the accuracy gain of uniform re-asking (UR).
-This overhead is far lower than techniques like self-consistency, which might use 5–10 calls per query (a 500%–1000% overhead) to sample multiple reasoning paths for every question [9, 41].
-It is also cheaper than strategies that always do chain-of-thought reasoning for every question.
-Our method is thus relatively lightweight.
-In interactive applications, this could translate to slightly longer latency only on the hard queries where the model was unsure, which might be a reasonable trade-off.
-One could even adjust the entropy threshold on the fly depending on desired speed vs.
-accuracy: for example, in a time-critical setting, only trigger the second pass if entropy is extremely high (i.e.
-only when the model is very unsure), whereas in a high-accuracy-required setting, use a lower threshold to double-check more often.
-  • Comparison to “Step-Back” Prompting: Recently, some anecdotal reports and community forums have suggested manually prompting strategies where one asks an LLM to double-check an answer or take a “step back” when uncertain.
-Our approach formalizes and quantifies this idea by tying it to an actual uncertainty metric.
-Rather than always asking “Are you sure?” after every answer (which would be tedious and could even degrade performance if done when not needed), we selectively apply a step-back prompt when the model’s own probabilities indicate uncertainty.
-This selective trigger is a key novelty – it bridges confidence estimation with adaptive prompting.
-Essentially, we give the model a chance to reflect only when it hesitated in the first place.
-This focus makes our method both efficient and targeted, avoiding unnecessary second-guessing when the model was already confident (and likely correct).
+  • Generality: Although we focus on multiple-choice QA, the idea can extend to open-ended tasks by triggering re-asking when uncertainty is high (e.g., low likelihood of the top completion or high entropy over continuations).
+This is harder without discrete options, but prior work defines entropy in open generation by clustering semantically equivalent answers [22].
+A practical variant is: answer once, estimate confidence (e.g., log-likelihood or sampling variance), then request justification and revision when confidence is low.
+  • Efficiency Considerations: Our method adds at most one extra call on flagged queries.
+In targeted re-asking (TR), this is 7,347 extra calls for 12,032 queries (61% more total calls) while achieving 99.6% of the gain from uniform re-asking (UR).
+This is far cheaper than self-consistency (5–10 calls per query; 500%–1000% overhead) [10, 23] or always-on CoT.
+In interactive settings, latency increases mainly on hard, uncertain queries, and thresholds can be tuned for speed–accuracy trade-offs.
+  • Comparison to “Step-Back” Prompting: “Step-back” prompts ask models to double-check, but applying them universally can waste compute and even hurt performance.
+Our method formalizes this idea by tying it to an uncertainty metric: we trigger reflection only when the model’s probabilities indicate hesitation.
+This selective trigger is central to the efficiency and avoids unnecessary second-guessing on confident (often correct) cases.
 VIII Limitations
-Despite the encouraging results, our study has clear limitations.
-At the same time, they point to concrete next steps: strengthening uncertainty calibration across model families, extending entropy-triggered deliberation beyond multiple-choice tasks, and making the second-pass prompting policy more robust while remaining compute-aware.
+Despite the encouraging results, our study has clear limitations that suggest next steps: improving uncertainty calibration across model families, extending entropy-triggered deliberation beyond multiple-choice tasks, and making the second-pass prompting policy more robust while remaining compute-aware.
 We discuss these limitations below.
-  • Dependency on Model Calibration: A primary limitation of our approach is that its effectiveness is not uniform across architectures and appears to be deeply intertwined with a model’s training history, particularly concerning potential data contamination.
-We hypothesize that the minimal accuracy gains observed in the Qwen3 series, despite high re-asking rates, might indicate prior exposure to the MMLU-Pro dataset during its extensive post-training or alignment phases.
-In contrast, both the DeepSeek V3.2 and V3.2 Reasoner models demonstrated substantial performance improvements.
-This suggests these models likely had not been exposed to similar benchmark data during their training.
-Therefore, the observed accuracy gains can be more confidently attributed to the Reinforcement Inference process itself, which successfully guided the models to correct their reasoning in areas of genuine uncertainty.
-This positions our method as a tool for enhancing real-time reasoning rather than merely recovering memorized information.This highlights a critical challenge for generalization: the utility of our method is diminished on benchmarks that a model may have already been trained on.
-Its success is contingent on a model’s capacity for genuine generalization.
-Consequently, deploying this framework requires careful consideration of a model’s training provenance to ensure it can be meaningfully applied.
-  • Multiple-Choice Focus: The current implementation is tailored to multiple-choice questions, where evaluating correctness is straightforward and the model’s output distribution is discrete over the given options.
-For open-ended tasks (free-form QA, code generation, etc.), applying a similar strategy requires more nuance in defining uncertainty and perhaps generating an alternative solution.
-It is doable – for example, one could sample multiple answers or use dropout to get an ensemble of outputs, and compute a form of entropy – but we did not explore that here.
-Thus, our results directly speak only to tasks with a closed set of answer options.
-Additional research is needed to generalize the approach to open-ended tasks.
-  • Prompt Specificity: We used one particular phrasing for the second-pass prompt and observed good results, but we did not exhaustively compare alternatives.
-It’s possible that certain wordings could further improve effectiveness, or conversely that our prompt might not generalize to all contexts.
-For example, explicitly instructing the model “explain your reasoning before giving a final answer” in the second pass might yield even higher gains by ensuring it works through the problem (at the cost of longer outputs).
-We opted for a concise prompt to keep things simple.
-Future work could experiment with variations: there is a balance between nudging the model and over-constraining it, and we mostly gave it a light nudge.
-We also did not try adding few-shot examples of corrected reasoning; doing so might improve performance but would move away from our zero-shot paradigm.
-  • Anchoring and prompt ablations: Our second-pass prompt includes the model’s own first-pass output (“Your previous answer was: …”) and explicitly states that the last output was uncertain (high entropy).
-While this still reveals no ground-truth label, it can introduce anchoring effects (the model may implicitly defend or over-correct relative to its prior output).
-We have included a prompt-only re-asking ablation (Section VI-D) showing that generic “think step-by-step” encouragement alone does not explain the gains; nevertheless, a more complete study should further ablate removing the prior output and/or removing the entropy cue to isolate each component.
-  • No Adversarial or Trick Questions: Our evaluation did not include questions deliberately designed to fool the model’s uncertainty mechanism – for example, questions with a very appealing incorrect option that the model is 100% confident in (due to a common misconception), or cases where any hint of uncertainty leads the model to an incorrect line of reasoning.
-In such adversarial cases, our method would not trigger at all (because the model would be overconfident and not realize it’s wrong), or it might trigger but still fail to correct the error if the model’s reasoning is fundamentally flawed.
-We did not test on adversarial data, so our results reflect performance on normal challenging questions.
-Techniques to handle overconfident mistakes – perhaps by integrating an external verifier or by training the model to better recognize when it could be wrong – would be needed to address these scenarios.
-Essentially, Reinforcement Inference can only help when the model has some internal doubt or ambivalence; if the model is confidently wrong, other remedies are required (and identifying such cases is an open problem).
+  • Dependency on Model Calibration: Effectiveness is not uniform across architectures and may depend on multiple factors, including training provenance (e.g., possible benchmark exposure), model architecture, and baseline performance ceilings.
+For example, the minimal gains on Qwen3 despite high re-asking rates could reflect prior exposure to MMLU-Pro during post-training/alignment, but could also indicate that the model is already strong (or more decisive) on this benchmark, leaving limited marginal room for improvement.
+In contrast, DeepSeek-v3.2 and v3.2-Reasoner show substantial improvements.
+Overall, Reinforcement Inference is most useful when the model exhibits genuine uncertainty and there is headroom for correction; deployments should therefore consider both training history and baseline accuracy when selecting benchmarks and operating points.
+  • Multiple-Choice Focus: Our implementation targets multiple-choice QA, where correctness is clear and uncertainty is naturally defined over a discrete option set.
+Extending to open-ended tasks (free-form QA, code generation) requires more careful uncertainty definitions and alternative-solution generation (e.g., sampling/ensembles to estimate entropy), which we do not study here.
+  • Prompt Specificity: We adopt a single fixed second-pass prompt and do not systematically explore alternative phrasings.
+While different wordings may yield further improvements, they could also increase verbosity or introduce additional tuning effects.
+Moreover, incorporating few-shot corrective examples may enhance performance, but would depart from the zero-shot setting we aim to preserve for comparability and simplicity.
+  • Anchoring and prompt ablations: The second-pass prompt includes the first-pass output (“Your previous answer was: …”) and an explicit entropy cue, which may introduce anchoring (defending or over-correcting relative to the prior answer).
+While our prompt-only ablation (Section VI-D) shows generic CoT encouragement alone does not explain gains, future work should further ablate removing the prior output and/or entropy cue to isolate each component.
+  • No Adversarial or Trick Questions: We do not test adversarial cases designed to exploit overconfidence or mislead uncertainty signals.
+In such settings, the method may not trigger (confidently wrong) or may still fail if reasoning is flawed.
+Addressing overconfident mistakes likely requires external verification or improved self-uncertainty training; identifying such cases remains open.
 IX Future Work
-The encouraging results from this initial study open up several directions for future work.
-In particular, we view Reinforcement Inference as an instance of a broader test-time control paradigm: using internal uncertainty estimates to allocate additional compute only when it is most likely to help.
-Realizing this vision robustly requires moving beyond a single benchmark and a single prompting template, and carefully characterizing when uncertainty signals are reliable versus misleading.
-The directions below therefore focus on generality, robustness, and system-level policies that make uncertainty-aware deliberation practical in real deployments:
-  • Broader Evaluation: We plan to evaluate Reinforcement Inference on a broader range of tasks and models.
-For instance, testing with other advanced models like GPT-4 or newer open-source models would reveal whether they benefit similarly from uncertainty-triggered reasoning.
-This may require creative approaches if raw probabilities are not directly accessible (for some API models, one might prompt the model to estimate its confidence [11], although models can misreport confidence if not calibrated).
+Encouraging results from this initial study motivate several directions for future work. The directions below emphasize generality, robustness, and deployable system policies:
+  • Broader Evaluation: We plan to evaluate Reinforcement Inference across more tasks and models.
+Testing other advanced models and newer open-source systems would clarify whether uncertainty-triggered reasoning generalizes.
+When raw probabilities are unavailable, one may need alternatives such as prompting confidence estimates, though these can be unreliable without calibration.
 Expanding to more benchmarks (e.g.
-other MMLU categories, reasoning-heavy datasets like GSM8K for math, or coding challenges) will help verify the generality of the method.
-  • Dynamic Thresholds and Policies: In a deployed system, one could dynamically adjust the entropy threshold or use more sophisticated triggering policies.
-For example, a system could start with a certain $\tau$ but increase it if too many second passes are being triggered with little benefit (to save time), or lower it if the goal is maximum accuracy and extra compute is acceptable.
-One could even train a small classifier to predict when a second pass would likely be beneficial, based on features of the probability distribution and question.
-Automating the decision of when to invoke extra reasoning – essentially learning a meta-policy – is an interesting avenue.
-  • Multi-Step Reinforcement: We limited our method to one extra pass.
-But what if the model is still uncertain after the second pass?
-One could consider iterating further: for example, a third pass with a prompt like “You are still unsure, try one more time considering each option carefully.” There may be diminishing returns, but a small number of iterations could potentially convert a few more hard cases.
-We would need to be cautious to not enter an endless loop; setting a maximum of perhaps two iterations beyond the initial attempt might be reasonable.
-Investigating how often a second pass remains uncertain (and if a third could help) is worthwhile future work.
-  • Integrating Verification or Refusal: In critical applications, one might combine Reinforcement Inference with a final verification stage.
-For instance, if the model remains uncertain (high entropy) even after the reinforced reasoning attempt, perhaps it should then abstain or escalate to a human.
-In that sense, our approach could be part of a larger pipeline: (i) model gives an initial answer, (ii) if unsure, model tries a reinforced answer, (iii) if still unsure or the answer is suspect, the system refuses or hands off to a human.
-This would maximize the model’s chances to answer correctly on its own, but still maintain safety when the model truly doesn’t know the answer.
-  • Applications in Generative Tasks: While we focused on Q&A, the idea of using uncertainty to trigger a self-correction could apply to generative tasks like story generation or code synthesis.
-For example, if a model is writing code and its internal probability distribution shows high entropy about the next steps (perhaps indicating confusion about how to proceed or a potential error), one could prompt it to “double-check the last steps” or generate an alternate solution.
-Similarly, in story or essay generation, if the model is uncertain, it might benefit from a prompt to rethink coherence or consistency.
-Exploring how to detect and act on uncertainty in creative generation is an intriguing direction.
-  • Extending Reinforcement Inference to VLA and Robotics: An exciting direction for future work is to extend Reinforcement Inference beyond discrete answer selection to vision–language–action (VLA) and robotic control models [48, 47].
-Modern VLA systems are increasingly decoder-based and autoregressive, generating action sequences token by token, which makes them a natural fit for uncertainty-triggered refinement.
-    In this setting, entropy can be defined over the action distribution (e.g., discretized action tokens, spatial bins, or latent action representations), providing a measure of execution uncertainty.
-High spatial or action entropy corresponds to diffuse, unstable action proposals, while low entropy indicates confident, convergent trajectories.
-Reinforcement Inference could then be used to trigger additional refinement steps when the model remains uncertain, encouraging actions to progressively concentrate in high-certainty regions of the action space.
-    This perspective aligns closely with human motor behavior: when performing a task such as grasping an object, initial movements are exploratory and uncertain, but as the hand approaches the target, actions naturally become more precise and less ambiguous.
-By explicitly conditioning additional action-generation passes on uncertainty signals, a VLA model may similarly improve reliability without relying on external rollouts, value functions, or explicit planning.
-    Unlike traditional imitation learning, which optimizes for average behavior [50], or reinforcement learning, which depends on task-specific rewards, an entropy-constrained autoregressive approach focuses directly on reducing execution uncertainty.
-This idea is also compatible with modern visuomotor policy learning approaches (e.g., action diffusion) [49].
-Exploring how such uncertainty-aware action refinement interacts with existing VLA and robotic learning paradigms is a promising direction for future research.
-  • Entropy-Guided Training: Our work hints that models implicitly know when they’re unsure, but they are not optimized to act on that knowledge by default.
-A future direction is to incorporate uncertainty signals into the training process itself.
-For instance, one could add a auxiliary loss term that encourages the model to produce low entropy (high confidence) distributions only when it is correct, and high entropy when it is truly unsure.
-This would push the model to be confidently correct and appropriately hesitant when it doesn’t know, improving calibration.
-Recent research has stressed that LLMs must be taught to know what they don’t know [51].
-Instead of only fine-tuning for better calibration post hoc [13, 28], one could train a model from the ground up with objectives that penalize incorrect high-confidence predictions.
-If successful, such a model would not only make fewer mistakes, but also be a perfect candidate for reinforcement inference techniques – its internal uncertainty signals would be even more trustworthy.
-In the long run, this line of research could lead to models that are jointly optimized for accuracy and for the correctness of their own confidence estimates.
+other MMLU categories, reasoning-heavy datasets like GSM8K, or coding challenges) will further test generality.
+  • Dynamic Thresholds and Policies: In deployment, thresholds (e.g., $\tau$) need not be fixed: a system could raise $\tau$ when second passes yield little benefit or lower it when maximizing accuracy is prioritized.
+More generally, one could learn a lightweight meta-policy that predicts whether a second pass will help from features of the probability distribution and question.
+  • Multi-Step Reinforcement and Verification: We currently allow only one extra pass.
+If the model remains uncertain after the second attempt, one could allow a small number of additional iterations while enforcing a strict cap to avoid loops.
+In critical applications, this can be paired with verification or abstention: (i) produce an initial answer, (ii) re-ask if uncertain, and (iii) if uncertainty remains high after reinforcement, refuse or escalate to a human.
+Measuring how often uncertainty persists after the second pass—and whether additional iterations or verification helps—is an open question.
+  • Applications in Generative Tasks: Although we focused on Q&A, uncertainty-triggered self-correction may apply to generative settings such as code synthesis or long-form writing.
+High-entropy regions during generation could trigger a “double-check” or alternative-solution attempt to improve correctness or coherence.
+Understanding how to detect and act on uncertainty in open-ended generation is an intriguing direction.
+  • Extending Reinforcement Inference to VLA and Robotics: A natural extension is to move from discrete answer selection to vision–language–action (VLA) and robotic control models [24], which are increasingly decoder-based and autoregressive and thus admit uncertainty-triggered refinement over action sequences.
+Entropy over the action distribution can serve as an execution-uncertainty signal: high entropy corresponds to diffuse, unstable action proposals, while low entropy indicates convergent trajectories.
+Reinforcement Inference could then allocate extra refinement only when this uncertainty is high, encouraging trajectories to concentrate toward high-certainty regions—a pattern that parallels human motor control, where early motion is exploratory and later motion becomes precise.
+Compared to imitation learning [25] or reward-driven reinforcement learning, this entropy-constrained autoregressive perspective directly targets reducing execution uncertainty and remains compatible with modern visuomotor policy learning methods [26].
+  • Entropy-Guided Training: Our results suggest models often know when they are unsure, but are not trained to use that signal by default.
+A natural direction is to integrate uncertainty into training, e.g., adding an auxiliary objective that encourages low entropy (high confidence) only when correct and higher entropy when genuinely uncertain, improving calibration.
+Recent work emphasizes that LLMs must learn what they do not know [27].
+Rather than relying solely on post-hoc calibration fine-tuning [12, 5], one could train with penalties for incorrect high-confidence predictions from the outset, yielding models that are both more reliable and better suited for reinforcement inference due to more trustworthy uncertainty signals.
 X Conclusion
-We introduced Reinforcement Inference, an inference-time approach that improves an LLM’s answers by leveraging its own uncertainty signals.
-Through experiments on 12,032 challenging multiple-choice questions from MMLU-Pro across 14 subjects, we demonstrated that when DeepSeek-v3.2 exhibited high predictive entropy—a proxy for hesitation—that uncertainty often signaled an incorrect first-pass answer, and that a second, more deliberate attempt could correct a large fraction of these errors without any retraining.
-This work highlights several key insights.
-First, large language models like DeepSeek do have measurable internal confidence; in our tests at scale, wrong answers were usually accompanied by higher entropy and lower max probability than correct answers.
-This underscores the importance of developing techniques to read and use those internal signals for more reliable AI behavior.
-Second, a simple second-pass prompt (essentially telling the model, ”you were unsure, so be careful now”) can unlock better reasoning in the model.
-The success of this prompt shows how far prompt-based control can go in steering model cognition–in this case, nudging the model from a fast guess toward a more deliberative mode.
-Third, our method achieved substantial gains at scale: overall accuracy rose from 60.72% to 84.03% (+23.31%) while re-asking 61.06% of questions.
-A uniform re-asking ablation (UR) in which we re-asked 100% of questions yielded only +0.10% additional accuracy over targeted re-asking (TR), indicating that uncertainty-based triggering is the key–not the volume of re-asking–and that uncertainty is central to addressing hallucination when the LLM is confused.
-Nearly identical results with 39% less computation demonstrate both the efficiency and the correctness of our design.
-Crucially, all of this was done at inference time, preserving the model’s weights.
-This has practical implications: even when one cannot retrain a model (due to cost, or if using a fixed API model), it may still be possible to improve its reliability through intelligent post-processing or on-the-fly prompting of its outputs.
-Reinforcement Inference contributes to the evolving toolbox for aligning and controlling LLM outputs [34].
-It resonates with the idea of making models not just give an answer, but also reflect on how they arrived at it and whether they are satisfied with it.
-By operationalizing a form of self-reflection triggered by uncertainty, we move a step closer to LLMs that can know what they know–and know what they don’t.
-We hope this work inspires further research into uncertainty-aware AI and meta-cognitive prompting techniques, which together could make AI systems not only more accurate, but also more self-aware and trustworthy in their reasoning.
-More broadly, our method is tightly coupled to the autoregressive nature of modern decoder-based models: because answers (and more generally, sequences) are generated token by token from a learned conditional distribution, entropy and related confidence measures arise naturally as first-class control signals during generation.
-Reinforcement Inference exploits this coupling by turning distributional uncertainty into an adaptive inference policy, allocating an additional pass only when the model’s own posterior indicates hesitation, thereby converting internal ambiguity into measurable and actionable compute.
-The same uncertainty-triggered refinement principle naturally extends beyond discrete answer selection to sequence-generation settings, including vision–language–action (VLA) models and robotic control.
-In such systems, spatial or action entropy can be interpreted as a proxy for execution uncertainty: high entropy corresponds to diffuse, unstable action proposals, whereas low entropy indicates confident, convergent trajectories.
-This perspective closely parallels human motor behavior, where early movements are exploratory and imprecise, progressively concentrating into precise actions as the hand approaches the target, motivating uncertainty-aware action refinement as a foundation for improving reliability in autoregressive VLA generation.
-Stepping back, the broader message is that uncertainty is not merely a diagnostic; it is a control signal.
-Treating entropy as a first-class input to an adaptive inference policy provides a principled way to allocate compute, probe a model’s latent capability horizon, and improve reliability without changing model weights.
+We introduced Reinforcement Inference, an inference-time method that leverages a model’s uncertainty to selectively trigger a second, more deliberate attempt.
+On 12,032 MMLU-Pro multiple-choice questions across 14 subjects, DeepSeek-v3.2 improves from 60.72% to 84.03% (+23.31%) while re-asking 61.06% of questions; re-asking 100% yields only +0.10% additional accuracy over targeted re-asking (TR), indicating that uncertainty-based triggering captures most of the gain with substantially less compute.
+More broadly, our framework is naturally aligned with the autoregressive nature of modern decoder-based models: because outputs are generated token by token from a learned conditional distribution, entropy and related confidence measures arise directly during generation and can be used as control signals.
+Reinforcement Inference exploits this coupling by turning distributional uncertainty into an adaptive inference policy—allocating extra compute only when the model hesitates—to improve reliability without changing model weights.
+This perspective complements broader efforts to align and control LLM behavior and motivates further work on uncertainty-aware inference and training.